### v1

# Agnostics: Learning to Code in Any Programming Language via Reinforcement with a Universal Learning Environment

Abstract
Large language models (LLMs) already excel at writing code in high-resource languages such as Python and JavaScript, yet stumble on low-resource languages that remain essential to science and engineering. Besides the obvious shortage of pre-training data, post-training itself is a bottleneck: every new language seems to require new datasets, test harnesses, and reinforcement-learning (RL) infrastructure.
We introduce Agnostics, a language-agnostic post-training pipeline that eliminates this per-language engineering. The key idea is to judge code solely by its externally observable behavior, so a single verifier can test solutions written in any language. Concretely, we (i) use an LLM to rewrite existing unit-test datasets into an I/O format, (ii) supply a short configuration that tells the verifier how to compile and run a target language, and (iii) apply reinforcement learning with verifiable rewards (RLVR) in a robust code execution environment.
Applied to five low-resource languages—Lua, Julia, R, OCaml, and Fortran—Agnostics (1) improves Qwen-3 4B to performance that rivals other 16B–70B open-weight models; (2) scales cleanly to larger and diverse model families (Qwen-3 8B, DeepSeek Coder 6.7B Instruct, Phi 4 Mini); and (3) for ${\leq}16$B parameter models, sets new state-of-the-art pass@1 results on MultiPL-E and a new multi-language version LiveCodeBench that we introduce.
We will release the language-agnostic training datasets (Ag-MBPP-X, Ag-Codeforces-X, Ag-LiveCodeBench-X), training code, and ready-to-use configurations, making RL post-training in any programming language as simple as editing a short YAML file.
1 Introduction
Figure 1: Overview: Agnostics Data Preparation and Training. (1) We reformulate existing coding datasets to our format. (2) We adapt the language-agnostic datasets to a particular programming language. (3, 4) We reinforce coding via Group Relative Policy Optimization (Shao et al. 2024; DeepSeek-AI et al. 2025), verifying the programs in our code execution sandbox.
Large language models (LLMs) are remarkably good at programming tasks, especially when codinng in high-resource programming languages such as Python and JavaScript. Their proficiency in low-resource programming languages, such as Fortran, Julia, and others, is significantly more limited. This gap appears both on benchmarks (Cassano et al. 2023) and in popular discourse. However, many low-resource languages are widely used in computational science (e.g., Fortran), medicine (e.g., Mumps), engineering (e.g., Matlab), and others sectors. For programmers in these sectors to be able to take full advantage of LLMs, we need methods to make LLMs better at the languages that they use.
The capability gap between high-resource and low-resource languages occurs for two reasons. The first one is that there is vastly more training data for some programming languages. For example, The Stack V2 (Lozhkov et al. 2024a), the largest public training corpus of code, has ${\approx}200$GB of Python but only ${\approx}2$GB of Julia and Fortran. Thus pretraining on code makes models significantly better at Python. A more subtle reason is the availability of post-training datasets and techniques. Contemporary LLMs are developed with an extensive post-training process that relies on (a) high-quality curated data for supervised fine-tuning, and (b) carefully designed environments for reinforcement learning, which must be able to execute and verify model-generated solutions. Both of these require significant human expertise, which is hard to find for low-resource languages.
This paper aims to dramatically facilitate post-training LLMs on low-resource programming languages, working towards closing the resource gap. Our key idea is that for a significant class of programming tasks, correctness can be specified as a property not of individual functions or code snippets, but of the entire program’s observable behavior (e.g., I/O). Furthermore, if its correctness can be tested with another program, we can use such a verifier together with appropriate problems and test cases to make a universal reinforcement learning environment which can be adapted to any programming language. In fact, the language used to implement the verifier and the one being learned are entirely independent. This approach matches how some existing post-training datasets are formulated (even though they are intended for Python/C++). We can also use LLMs to reformulate language-specific datasets into this format.
Based on this insight, our approach, Agnostics, works as follows (Figure 1). 1) We use an LLM to reformulate language-specific datasets into our standard language-agnostic format. 2) We generate prompts and a configured verifier targeting a particular language, with a small (4-5 line) configuration. 3) We apply reinforcement learning with verified rewards (RLVR) using a robust, language-agnostic execution sandbox that we develop. 4) The result is a model specialized to the target language. Agnostics particularly excels at finetuning models for low-resource languages, as it does not rely on high-quality datasets specific to a particular language.
Contributions
We make the following contributions.
  1. 1.
    Agnostics, a post-training pipeline for coding in arbitrary programming languages, including low-resource ones;
  2. 2.
    The best-performing open-weights models for Lua, R, Julia, OCaml and Fortran at the ${\leq}16$B parameter scale;
  3. 3.
    3 Agnostics datasets Ag-MBPP-X, Ag-Codeforces-X, and Ag-LiveCodeBench-X, based on MBPP (Austin et al. 2021), Open-R1 Codeforces (Penedo et al. 2025) and LiveCodeBench (Jain et al. 2024a) respectively.
  4. 4.
    A small and highly-optimized Agnostics training framework, implementing GRPO and including a parallel code execution sandbox, sampling, rewards computation, and model back-propagation.
Artifacts and code necessary to verify the results we present are listed at agnostics.abgru.me. The training framework will be publicly released together with the formal publication of this work.
2 Background and Related Work
[…]
Figure 2: High-resource languages (left) dominate the Stack v2. We are interested in low-resource languages (right).
Data Scarcity for Low-Resource Languages
LLM have been pretraining on code for several years, both because LLMs are widely used for practical programming tasks, and because pretraining on code improves their general reasoning abilities (Ma et al. 2023). In addition to general-purpose LLMs, there are also code-specialized models that are either trained exclusively on code (e.g., Xu et al. (2022); Allal et al. (2023)) or trained on code starting from a general-purpose model checkpoint (e.g., Rozière et al. (2024)).
However, the publicly available pretraining data for code is heavily skewed toward a handful of programming languages. For example, consider The Stack V2 (Lozhkov et al. 2024b), which is the largest public code pretraining dataset, with code from GitHub and dozens of other sources. The Stack V2 dominated by a handful of programming languages: just 10 of 619 languages account for over 90% of the dataset. We are interested in developing models for low-resource languages that each account for less than 0.5% of the publicly available code (figure 2).
We can imagine working around this data scarcity in a few ways. However, previous work shows that up-sampling low-resource languages during pretraining only leads to small benchmark improvements (Orlanski et al. 2023), while fine-tuning on the pre-training data for low-resource languages has negligible impact (Cassano et al. 2024). Thus, it is not clear how to make further gains from existing natural data.
Synthetic Data for Low-Resource Languages
If natural data is not available for a task, it is possible to use LLMs to generate synthetic fine-tuning data (Wang et al. 2023), and there are many techniques for building code-centric supervised fine-tuning datasets (e.g. Luo et al. (2023); Wei et al. (2024c, b)) that work remarkably for Python. Although these approaches could in principle be applied to any programming language, Cassano et al. (2024) show that without distillation or verification (e.g., see Hu et al. (2025); Wei et al. (2024a)), synthetic tasks and code for low-resource programming languages are low-quality, and models fine-tuned on them perform poorly.
MultiPL-T (Cassano et al. 2024), similar to TransCoder-ST (Roziere et al. 2021) and CMTrans (Xie et al. 2024), couples synthetic data generation with verification using rejection sampling: it generates up $n$ candidate programs in a target, low-resource language and only fine-tune models on generations that pass hidden unit tests that it translates from Python. However, the MultiPL-T approach has two significant limitations. (1) For the verifier to not reject all samples, the model must be able to generate a working program within $n$ attempts. In MultiPL-T, ${\approx}30\%$ of prompts produce a working program for $n\in\{50,100\}$ attempts, and the rest are discarded. We train on much harder problems, and estimate that rejection sampling would require an order of magnitude more resources to achieve a comparable acceptance rate (§ A). (2) For each low-resource language of interest, MultiPL-T requires the user to write a little compiler to translate test cases and function signatures from Python to the target language. MultiPL-T only supports a limited set of built-in Python types (e.g., no classes) and it dictates that all Python types and values must faithfully map to the target language. However, depending on the problem and language, the natural way to represent data may not map cleanly to Python types. This can lead to peculiar, unidiomatic translations that require deep language expertise to get right. The Agnostics approach is significantly easier to use than MultiPL-T, and only requires the user to know how to compile and run a program in the target language from the shell.
Reinforcement Learning on Coding Tasks
DeepSeek R1 (DeepSeek-AI et al. 2025) popularized RL on LLMs with rule-based rewards, instead of learned reward models. R1 reports applying RL to coding tasks but does not disclose further dataset details. A number of papers apply RL to the NL to code task (Zeng et al. 2025; Gehring et al. 2024; Jain et al. 2025). These techniques, which target Python, show that reinforcement learning can improve LLM capabilities beyond what is possible with just supervised fine-tuning.
However, the key benefit of reinforcement learning is that it can train a model to do tasks for which high-quality data for supervised fine-tuning is unavailable. There are recent examples of using RL for code optimization (Du et al. 2025; Nichols et al. 2024), resolve GitHub, issue resolution (Wei et al. 2025), and iterative development (Zhou et al. 2025). These papers target tasks in high-resource languages—C++, Java, and Python—whereas Agnostics targets several low-resource languages.
3 The Agnostics Approach
Our approach comprises (1) a data preparation stage which reformulates language-specific programming tasks to be language-agnostic, and retargets language-agnostic datasets to a programming language of interest (1, 2 in figure 1); and (2) the training stage which uses the GRPO algorithm with an efficient, language-agnostic verification framework (3, 4 in figure 1). Our central idea is to have tasks ask for programs which behave in a particular way. The test cases are samples of this behavior, and a verifier program can check if a solution behaves according to the sample. In this paper, we limited ourselves to working with tasks asking for programs which read data from the standard input, compute a unique answer, and write it to the standard output. Hence, the datasets we prepared share the same verifier.
# Write a python function to identify non-
# prime numbers.
def is_not_prime(n):
assert is_not_prime(2) == False
assert is_not_prime(10) == True
(a) A task prompt and associated tests (in gray) from MBPP.
Instruction: Given an integer NN ($N\geq 2$), determine whether it is a non-prime number. Output ‘True’ if the number is non-prime, ‘False’ otherwise. Input format: a single integer NN ($N\geq 2$). Output format: a single line containing ‘True’ or ‘False’.
[…]
(b) The task and tests reformulated for Agnostics.
Figure 3: For dataset preparation, we use an LLM to reformulate fine-tuning datasets with language-specific prompts and tests (above) into equivalent language-agnostic programming tasks.
3.1 Dataset Preparation
Some datasets, like Open-R1 Codeforces (Penedo et al. 2025), already define tasks in the desired I/O style. More commonly, however, code datasets provide a set of unit tests. Figure 3(a) shows a representative item from MBPP: it has a natural language problem description and a Python function signature that comprise the prompt, and a suite of tests used to test model-generated code. These datasets can be easily translated into the I/O format.
To make such problems language‑agnostic and compatible with our verifier, we prompt an LLM to reformulate each task so that the program communicates exclusively via plain‑text standard in and standard out. We ask the model to spell out concrete I/O conventions—number of decimal places, newline versus comma separators, ordering of values, and so on—so that the expected behavior is unambiguous. Figure 3(b) shows the reformulated example. § B has the instruction we use to reformulate MBPP; other datasets might require small changes to the prompt.
3.2 Programming Language Preparation
install: apt-get install -y r-cran-tidyverse
filename: snippet.R
execute: Rscript snippet.R
prompt: |
Use R version 4.
Use readLines(con = file("stdin"))‘ to
read input from stdin. Optionally, use
the n argument to read the first n
lines. For example:
‘‘‘r
input <- readLines(con = file("stdin"), n = 1)
n <- as.integer(input)
print(n) # print the first line of input
Also, use cat to print output to stdout.
For example:
‘‘‘r
cat(n)
Please do not use print to print output.
Figure 4: An Agnostics configuration snippet for R.
To prepare a new language, we author a small configuration file that serves two purposes. First, it defines a prompt prefix (prepended to each problem by the trainer) which instructs the model to produce code in the target language. Second, the configuration file specifies the shell commands to install the language toolchain and run code. Figure 4 shows the configuration for the R language, with an atypically large prompt due to quirks of its I/O APIs.111 There are three ways to run R, three I/O APIs, and only one portable way to read from standard in.
A good prefix should help avoid the most common errors, but determining which errors are common can be time-consuming. In our experience, if the model starts from ${\geq}5\%$ accuracy, then the prefix only needs to say that the solution must be in language $L$. (This varies with GRPO group size, see § 3.3.) On the other hand, when starting from near-zero accuracy, a prefix like in figure 4 can help prevent common mistakes. For OCaml and Fortran, we asked a capable LLM (OpenAI o3) for advice based on several faulty generations.
What follows are several Fortran programs. You’ll see that most of them are wrong. Read them carefully and identify the Fortran programming mistakes that I’m making. Ignore algorithmic mistakes, and focus on my misconceptions about Fortran. Come up with advice on how I should program Fortran correctly. Distill this advice into 10-20 sentences.
We used the resulting instructions verbatim (§ C), but only for training and not to evaluate models on OCaml or Fortran.
3.3 Trainer and Code Execution
The Agnostics trainer uses the Group-Relative Policy Optimization (GRPO) reinforcement learning algorithm (Shao et al. 2024), with verifiable rewards (DeepSeek-AI et al. 2025), and some common tweaks to further improve its efficiency (Yu et al. 2025). We couple the algorithm with a language-agnostic code execution framework that is designed to be robust and efficient.
Trainer
The trainer instantiates the GRPO algorithm as follows. Let $(x,\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K})\sim\mathcal{D}$ be a dataset of language-agnostics tasks, where $x$ is the task prompt and $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$ is the set of I/O examples. Let $L$ be a language configuration $L$ (e.g., figure 4). From the behavior policy $\pi_{\theta_{\text{old}}}$, we sample a group $G$ of candidate responses
$\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(y_{i}\mid L.\mathsf{prompt},x).$
We assign each candidate a reward $R_{i}$, with $R_{i}=1$ if the execution environment (described later) verifies that the extracted program’s behavior is consistent with the I/O examples $(\mathit{in}_{k},\mathit{out}_{k})$, and $R_{i}=0$ otherwise.
We turn group rewards into sequence-level advantages
$\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}(\{R_{j}\}_{j=1}^{G})}{\operatorname{std}(\{R_{j}\}_{j=1}^{G})}.$
We then update the policy with the objective
$\mathcal{L}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Bigl{(}r_{i,t}(\theta)\hat{A}_{i},c_{i,t}(\theta)\hat{A}_{i}\Bigr{)},$
where
$c_{i,t}(\theta)$ $=\text{clip}\bigl{(}r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\bigr{)},$
$r_{i,t}(\theta)$ $=\frac{\pi_{\theta}(y_{i,t}\mid L.\mathsf{prompt},x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid L.\mathsf{prompt},x,y_{i,<t})}.$
We omit the KL-divergence term, similar to Yu et al. (2025).
We considered and decided against a reward for a partially-correct answer. We tried to reward the model for code which runs without errors but produces wrong output or for code which only passes the public tests. In both cases the models were very likely to learn how to exploit the reward, e.g., by producing empty programs or by hard-coding the public tests (and claiming to produce a “draft answer”).
Code Execution
Our verifier is a language-agnostic code execution sandbox which (1) extracts a program from each candidate; (2) compiles it if needed; and (3) tests it on the I/O examples $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$.
To extract the code, we instruct the model to put it in a Markdown block, which all major instruction-tuned models do by default. Since we rely on the native format of the model, we do not need to train the model with a format reward. This guarantees that the increases in rewards that we see during training are genuine improvements and not merely the result of the model learning to format correctly.
For each language, we build and cache an OCI container using the configuration $L$. To build the container, we install the language compiler and runtime (the script $L.\texttt{install}$), and include a generic execution harness which runs and tests candidate programs. The execution harness runs continuously in the container, waiting to receive a triple with the candidate program, the set of input/output examples, and timeouts. The harness (1) writes the program to disk (to $L.\texttt{filename}$), (2) compiles it if needed ($L.\texttt{compile}$), (3) runs it on each received input ($L.\texttt{execute}$), and verified that it produces the expected output. The harness imposes timeouts on the compilation step and each execution, and returns reward 0 on any timeout or failed verification.
It is important to have timeouts for both compilation and execution. This prevents pathologies such as unbounded macro expansion in Julia (caught by the compile timeout) and infinite loops (caught by the execution timeout). Using containers also allows us to limit CPU, memory, and filesystem usage; no elevated privileges are granted to the generated program. Although the current datasets only specify tasks by standard I/O, the same sandbox can safely accommodate problems that read/write from network or disk.
A subtle resource limit that we impose is a limit on the size of output. Even with a short timeout such as $30$ seconds, a pathological candidate program can output tens of gigabytes of text. This can crash the verifier if it naïvely tries to read and store all output. Instead, the verifier maintains a fixed-size (5MB) read buffer and immediately kills programs which overflow it.
Overall, this design lets us keep a pool of warm containers for the duration of training: we find that spawning a fresh container is two orders of magnitude slower than re-using an existing one. In our experiments, a single training run may involve testing 150,000 programs, each on several I/O examples. Most of the generated programs are faulty and some behave badly, e.g., they either timeout, consume too much memory, or produce too much output. So containers do occasionally crash or need to be killed, and our execution environment handles this automatically.
Finally, to improve compile times, our execution environment mounts a RAM disk in each container. Compilation may be slow due to creating many intermediate files, and indeed some large C++ projects, e.g., Firefox, recommend using a RAM disk to speed up their builds (Firefox 2025).
Implementation
We implement the trainer and execution environment with Ray (Moritz et al. 2018), which facilitates multiprocessing and distributed computing. In particular, the framework allows distributing the training over a network of heterogeneous nodes, which allows running the trainer on a node specialized for GPU work and the execution environment on a node specialized for CPU work. Ray also lets us easily have two separate inter-communicating processes, one for group generation and another for verification and loss computation, simply by defining two actor classes. Running the two concurrently significantly speeds up training, as we found that they take a roughly comparable amount of time. The execution environment is also implemented as an actor, with one Python asyncio task per container. Using tasks instead of actors helps lower how much inter-process communication and data copying is needed.
Hyperparameters
For our training, we use the AdamW optimizer (Loshchilov and Hutter 2019) with a learning rate of $5\times 10^{-6}$ and a cosine decay schedule with a warmup of 0.1 epochs. We process 4 prompts in each batch, with group size 32 per prompt. We generate with temperature 0.7. We train hybrid reasoning models with reasoning disabled. (We still often observe generations with reasoning-like text, either before the answer or in comments.)
4 Evaluation
To evaluate our approach and codebase, we train and benchmark multiple models on 5 low-resource programming languages. We measure pass@1 accuracy with reasoning disabled and 20 samples per prompt at temperature 0.2.
4.1 Training datasets
Ag-Codeforces-X, the main dataset we use for training, was created based on competitive programming problems from the Open-R1 Codeforces dataset (Penedo et al. 2025). Few adjustments were necessary, since the problems already specified programs and tests using standard I/O. The train split contains 5369 problems. See § B for more details.
4.2 Benchmarks
We evaluate Agnostics with the following benchmarks.
MultiPL-E (Cassano et al. 2023) is a well-established benchmark which has been frequently used to evaluate the performance of new LLMs on a broad set of languages (e.g., Kimi Team (2025); Yang et al. (2025); Grattafiori et al. (2024); ByteDance et al. (2025)). MultiPL-E was prepared by compiling HumanEval (Chen et al. 2021) prompts and unit tests from Python to each target language. Each MultiPL-E programming language requires a $\approx 500$ LOC prompt and test translator, considerably more effort than writing an Agnostics configuration file.
A major limitation of MultiPL-E is being too easy for frontier models. With Python, frontier models are now evaluated on solving programming contest problems (Jain et al. 2024a); no multi-language benchmarks are as challenging.
Ag-LiveCodeBench-X, a contribution of this paper, is a new multi-language benchmark derived from LiveCodeBench. LiveCodeBench 5.0 has 880 problems, of which 381 have Python starter code and test cases. The remaining 499 problems do not use starter code and instead use standard I/O to specify and test solutions. Hence we used these problems to transform LiveCodeBench into an Agnostics dataset. Accordingly, benchmarking a new programming language with Ag-LiveCodeBench-X is straightforward: we can reuse the language configurations and execution environment from our trainer (§ 3.3).
Importantly, the problems from LiveCodeBench and thus also Ag-LiveCodeBench-X are filtered out from our training sets. Moreover, our results show that Ag-LiveCodeBench-X is significantly harder than MultiPL-E.
4.3 Results
We now present our results. We use a few abbreviations in the tables. Ag-LCB-X stands for Ag-LiveCodeBench-X; we clarify abbreviated model names in the text. One table row may present scores of many models; the column for language X shows the score of the model trained on X.
Table 1: Pass@1 rates on AG-LiveCodeBench-X for Qwen 3 4B models trained on five low-resource languages using Ag-Codeforces-X. Each model achieves a SOTA result for an open-weight model with ${\leq}16$B parameters.
Model Ag-LiveCodeBench-X
X= Lua Julia R OCaml Fortran
Llama-3.3-70B-Ins 25 22 13 7 3
Qwen3-32B 22 26 17 2 1
DSCv2-Lite-Ins-16B 13 12 9 7 6
[…]
Table 2: Pass@1 rates on MultiPL-E for Qwen 3 4B models trained using Ag-Codeforces-X. The significant improvement indicates that they did not overfit to the competitive programming format.
Model MultiPL-E
X= Lua Julia R
Qwen3-4B 61 51 36
[…]
SOTA small LLMs for low-resource PLs
Using Agnostics, we train Qwen 3 4B on Ag-Codeforces-X specialized to Fortran, Julia, Lua, OCaml, and R. The resulting Qwen3-4B-CF-X models are, to the best of our knowledge, state-of-the art low-resource programming language code models with open-weights and ${\leq}16$B parameters.
Benchmarking them on Ag-LiveCodeBench-X (table 1), we see significant improvements. (i) On every language, the models match or outperform DeepSeek Coder v2 Lite Instruct (16B). (ii) On every language, our performance comes close to or even exceeds that of Qwen 3 32B and Llama 3.2 70B. (iii) Compared to the base model, Qwen 3 4B, pass@1 improves by a factor of 1.5–2x. It is safe to assume that the Qwen models, like the Llama models (Grattafiori et al. 2024), are trained on all the publicly available coding data; hence, Agnostics improves the models beyond what typical training on such data allows.
(iv) Finally, on OCaml and Fortran, we improve the performance of Qwen 3 4B from near zero to 7% and 15%. Notably, although we train the models with a prompt prefix to facilitate learning (§ 3.1), we do not use this prompt prefix for evaluation. Thus the pass@1 scores represent what the model learned, and not information provided in context.
A potential weakness of our approach is that our models are overspecialized to the competitive programming format, and can only generate programs that use standard I/O. We also evaluate with the established MultiPL-E benchmark, which does not use this format, and also find significant improvements (table 2). [^2] Note that MultiPL-E does not support OCaml and Fortran. This indicates that our models have not overfit to the competition programming format.
Figure 5: Training Qwen3-4B-CF-X, GRPO batch pass@1.
Figure 5 shows the GRPO batch pass@1 rates seen when training Qwen3-4B-CF-X. All the models follow similar curves, partially due to being trained on the same data permutation. Nearly all the models slowly keep improving almost until the dataset end. We also observed the train and test split rewards to be correlated with each other (§ D.2).
Table 3: Pass@1 rates for larger Qwen 3 8B models trained on three languages using Ag-Codeforces-X. Agnostics scales to models of this size.
Model MultiPL-E Ag-LCB-X
X= Lua Julia R X= Lua Julia R
Qwen3-8B 63 53 44 11 9 9
[…]
Agnostics scales to larger models
To test if improvements from Agnostics training scale with model size, we train the Qwen 3 8B model on Ag-Codeforces-X specialized to Lua, Julia and R and benchmarked it on Ag-LiveCodeBench-X and MultiPL-E (Table 3). The Qwen3-8B-CF-X models improve significantly on both benchmarks, improving over their 4B counterparts.
We expect Agnostics to scale to even larger models, with appropriate computing resources. However, we found that Agnostics training with Ag-Codeforces-X does not improve two smaller models: Qwen 3 1.7B and Llama 3.2 3B Instruct, perhaps due to the problems being too difficult.
Table 4: Pass@1 rates on for Qwen 3 4B models trained using MBPP. This smaller and easier dataset still results in improvements, though smaller than those in Tables 1 and 2.
Model MultiPL-E Ag-LCB-X
X= Lua Julia R X= Lua Julia R
Qwen3-4B 51 61 36 11 10 10
[…]
Agnostics works with easier problems
So far, all of the models that we have trained so far have used the Open-R1 Codeforces dataset. To show that Agnostics works with other datasets, we also train models for Julia, Lua, and R using the MBPP training set. (§ 3.1 describes how we prepare MBPP.) Note that the MBPP problems are trivial—see figure 3(a)—compared to the Open-R1 Codeforces problems. So, we cannot expect these models trained on the MBPP problems to be as good as the models we presented before.
Nevertheless, training on MBPP still improves performance on Lua and Julia (table 4). On R, the table shows a small drop in performance on Ag-LiveCodeBench-X, but an improvement on MultiPL-E.
Table 5: Pass@1 rates for DeepSeek Coder V1 and Phi 4 models trained using Ag-Codeforces-X. We find that Agnostics works on other model families.
Model MultiPL-E Ag-LCB-X
X= Lua Julia R X= Lua Julia R
Phi4-mini-ins 40 39 34 7 7 0
DSC-6.7B-Ins 40 54 37 8 5 7
[…]
Agnostics works on other model families
Finally, to test if Agnostics works on other models, we train both DeepSeek Coder 6.7B Instruct (Guo et al. 2024) and Phi 4 Mini Instruct (Microsoft et al. 2025) with Ag-Codeforces-X specialized to Julia, Lua, and R. Agnostics is able to improve the performance of both models on all languages, as measured by MultiPL-E and Ag-LiveCodeBench-X with one exception (Table 5). Phi 4 Mini Instruct scores zero on Ag-LiveCodeBench-R, and training does not improve the score. However, on the easier MultiPL-E benchmark, we see that the model is slightly better on simpler R problems ($34\rightarrow 35$).
Note that DeepSeek Coder 6.7B is a relative old LLM, superseded by the much larger DeepSeekV2 and V3 models. Unlike Qwen 3, DeepSeek Coder is not trained with reinforcement learning, but is only an instruction-tuned model. Thus this result also shows that Agnostics can work on models that have had relatively limited post-training.
Figure 6: We classify with an LLM what bugs appear in programs produced by Qwen 3 4B and our trained model, Qwen3-4B-CF-OCaml. Each radial represents a bug category, and points along the radial indicate how many programs appear to have that bug. Qwen 3 4B makes more fundamental programming mistakes, such as syntax errors and misusing builtins, indicating its limited grasp of OCaml. Our trained model makes significantly fewer such mistakes. We see a small increase in logic flaws: the trained model makes fewer shallow mistakes, revealing deeper issues.
4.4 Qualitative Improvements
In this section we take a deeper look at how training with Agnostics addresses the kinds of errors that Qwen 3 4B makes on low-resource languages. First, we define a taxonomy of common bugs by prompting an LLM (o3) to classify bugs in a sample of faulty programs and then lightly editing the suggestion. § E has the full taxonomy and the prompt we used to develop it. The taxonomy spans fundamental programming errors, such as syntax errors, and more subtle mistakes such as logic flaws.
We then sample 100 problems from Ag-LiveCodeBench-X, and take five programs produced by Qwen 3 4B and the Agnostics models trained on Open-R1 Codeforces problems. Recall from table 1 that these are significantly better than Qwen 3 4B. Using Sonnet 4, we use the taxonomy to classify the bugs in each program.
Figure 6 shows the bug distribution for OCaml. We see that the base Qwen 3 4B models makes substantially more fundamental programming mistakes in OCaml. More programs have syntax errors (55% vs 35% after training), more programs misuse builtin functions (60% vs 32%), and so on. However, we observe a small increase in logic flaws (18% vs 25% after training). Inspecting the results, the reason for this is that when a program is full of syntax errors and hallucinated functions, it is very difficult to even determine whether or not the algorithmic approach is correct. Training eliminates these shallow bugs and lets deeper issues manifest. We see the same patterns across the other four languages (§ E).
5 Conclusion
LLMs are strongest where pre-training and post-training data are abundant, and weakest where practitioners arguably need them most: for low-resource programming languages. This paper proposes Agnostics, a language-agnostic post-training pipeline that removes the per-language engineering tax by verifying code purely via externally observable behavior. A single verifier and a short configuration are enough to adapt the same reinforcement learning setup to new languages.
Empirically, Agnostics consistently improves small open-weight models on five low-resource languages–Lua, Julia, R, OCaml, and Fortran–without requiring language-specific test translators. Training Qwen-3 4B with Ag-Codeforces-X yields large gains on our new Ag-LiveCodeBench-X benchmark and on MultiPL-E, often rivaling or surpassing 16B–70B open-weight baselines. The method scales to larger and different model families: Qwen-3 8B benefits similarly, and we observe improvements on DeepSeek Coder 6.7B and Phi-4 Mini as well. Error-type analysis shows that the gains come from reducing fundamental language-usage mistakes.
A practical advantage of the approach is how little per-language work it requires. After the framework was in place, adding OCaml and Fortran took us less than an hour each. We expect adoption to be just as straightforward for any language with a command-line toolchain.
We see no reason to expect the approach to not work for models larger than 8B, although our experiments are limited by available compute. For scaling data, the Agnostics reformulation strategy apply to much larger problem sets as well. For instance, OpenCodeReasoning has $\sim$600K problems with Python solutions (Ahmad et al. 2025); converting such corpora into language-agnostic I/O tasks would provide rich RL datasets with many target languages with minimal additional engineering.
6 Acknowledgments
This work was supported by Czech Science Foundation Grant No. 23-07580X. This work is partially supported by the U.S. National Science Foundation (SES-2326174). This material is based upon work supported by the U.S. Department of Energy, Office of Science under Award Number DESC0025613.
We used GNU parallel (Tange 2023).
Disclaimer: This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
Appendix A Efficiency of Rejection Sampling
An alternative to reinforcement learning is to use rejection sampling with supervised fine-tuning: prompt the model to synthesize $n$ solutions to each task, reject solutions that fail tests, and fine-tune on the task-solution pairs that pass tests. If the model is weak at a task, which is the case with low-resource languages, then it is possible for all solutions to fail for a given task. Cassano et al. (2024) use this approach to get solutions for $\approx 30\%$ of the tasks in their dataset. They use $n=50$ for Lua and Julia, but increase to $n=100$ for OCaml.
The efficiency of this approach depends on the hardness of the task and the capabilities of the model. In this paper, we work with newer models that are marginally better at low-resource languages (based on MultiPL-E benchmark results). The task in Cassano et al. (2024) is to translate a simple, self-contained Python function from the model’s pretraining data into an equivalent function in another programming language. This task is significantly easier than the Agnostics task, which is to solve a competitive programming problem in a low-resource language, without any reference code or tests.
We can estimate the efficiency of rejection sampling from the rewards assigned in the early steps of reinforcement learning, when the policy is closest to the initial policy. Figure 7 shows the pass1 rates while training Qwen 3 4B on Fortran and OCaml. In the first 200 steps, the average pass rate for Fortran and OCaml are $1.9\%$ and $1.1\%$ respectively–considerably less than the $30\%$ success rate reported by Cassano et al. (2024) on their code translation task, making the rejection sampling approach prohibitively expensive for the low-resource language setting.
Appendix B Preparing Datasets for Agnostics
Ag-MBPP-X
is a dataset of Mostly Basic Programming Problems, transformed from the MBPP (Austin et al. 2021) dataset. The source dataset contains problems asking to complete a Python function from on its signature and a docstring, where the docstring specifies what result the function should return; our dataset contains equivalent programs which read/write the same data from/to standard input/output. Out of 974 problems, we are able to translate 776 problems into Ag-MBPP-X (348 problems are in the sanitized subset of MBPP).
Processing the dataset with Qwen3-32B took less than 1 hour using 2 H100 GPUs. Figures 3(a) and 3(b) in the main body of the paper shows a sample problem from the dataset, before and after the reformulation. We used the following prompt.
You are a competitive programming expert.
You are given a problem that asks you to implement a function.
Your task is to translate the description of the problem into a form that accepts one set of function arguments as inputs and return the function return value as output.
Use programming competition style input and outputs -- that is, priorize the use of spaces and newlines to separate inputs and outputs over using commas and parentheses (or other delimiters). Specifically, for 2d lists, you should print them as a list of lists, where the outer lists elements are separated by newlines and the elements of the inner lists are separated by spaces.
For example, a 2d list like [[1, 2], [3, 4]] should be printed as:
1 2
3 4
Do not use any other delimiters.
If there are multiple 2d lists, you should use 2 newlines to separate them.
for example, a 2d list like [[1, 2], [3, 4]] and [[5, 6], [7, 8]] should be printed as:
1 2
3 4
5 6
7 8
If the problem requires outputing decimal numbers, make sure the output format specifies to round all decimal numbers to 4 decimal places. In this case, you should also round all the numbers in the output to 4 decimal places.
Do not forget to specify the input and output format in the description.
Here is the problem description:
{original mbpp problem description}
Here are the test cases:
{original mbpp test cases}
You should return a json object with the following fields:
- "description": the description of the problem
- "input_format": a string describing the input format
- "output_format": a string describing the output format
- "tests": a list of test cases, each test case is a json object with the following fields:
- "input": a string that represents the input of the test case, in the same format as the input format in the description
- "output": a string that represents the output of the test case, in the same format as the output format in the description
Place your response in a single ‘‘‘json ‘‘‘ block. Do not include any other text in your response.
Ag-Codeforces-X
is a dataset of competitive programming problems, created from Codeforces problems in the open-r1/codeforces dataset. The source problems already specified programs by their I/O behavior, hence only very minor changes were needed to build language-universal Agnostics problems out of the fields in the dataset: we only skipped the time and memory restrictions present in the original problems. To be precise, we used data from the open-r1/codeforces-cots dataset, solutions_py_decontaminated subset, which contains problems decontaminated using 8-gram overlap against multiple benchmarks, in particular LiveCodeBench. We used the auxilliary checker_interactor subset to only keep the problems which admit a simple verifier for their solutions, i.e., problems where a single output is correct for each input and where the solution only needs to read data from its input, compute the result, and write it to the standard output. We prepared both a train and a test split. The former contains 5369 problems and the latter contains 107 problems we held out from the source dataset, 7 selected manually and 100 randomly. The train split features randomized prompts, which we found help with generalizing the results of training on the dataset to other benchmarks. The prompts were randomly split into a number of types. 30% of the prompts use standard Markdown headings to start different sections of the prompt, 35% use bold text instead, and the remaining 35% simply concatenate the prompt sections together. Half of the prompts of the final type also do not include an I/O sample in the prompt.
Ag-LiveCodeBench-X
is also a dataset of competitive programming problems, created from a subset of the LiveCodeBench dataset (Jain et al. 2024b). LiveCodeBench 5.0 has 880 problems, of which 381 have Python starter code and test cases. The remaining 499 problems do not use starter code and instead use standard I/O to specify and test solutions. Hence we used these problems to transform LiveCodeBench into an Agnostics dataset. Ag-LiveCodeBench-X only has a train split, like its source dataset.
Appendix C Agnostics Configurations
In this section, we list the configurations that we use for our target languages. The configuration files use YAML. The prompts for OCaml and Fortran have instructions generated by OpenAI o3.
The Lua configuration:
prompt: Use Lua 5.1, targeting LuaJIT.
install: apt-get install -y luajit
filename: snippet.lua
execute: luajit snippet.lua
The Julia configuration:
prompt: Use Julia 1.11.
container:
base-image: "julia:1.11.3"
type: debian
filename: snippet.jl
execute: julia snippet.jl
The R configuration (the same as figure 4, repeated here for convenience):
install: apt-get install -y r-cran-tidyverse
filename: snippet.R
execute: Rscript snippet.R
prompt: |
Use R version 4.
Use readLines(con = file("stdin"))‘ to
read input from stdin. Optionally, use
the n argument to read the first n
lines. For example:
‘‘‘r
input <- readLines(con = file("stdin"), n = 1)
n <- as.integer(input)
print(n) # print the first line of input
Also, use cat to print output to stdout.
For example:
‘‘‘r
cat(n)
Please do not use print to print output.
The OCaml configuration:
prompt: |
Use OCaml 5.
Numbers: + - * / mod vs. +. -. *. /. ** (add dots!)
Casts: float_of_int int_of_float int_of_string
Mutation: refs (:= !) or pass new values recursively
Strings: split_on_char, String.get => char, use Printf "%
Lists: avoid List.nth; prefer pattern-match / folds / arrays
container:
base-image: "docker.io/ocaml/opam:ubuntu-22.04-ocaml-5.0"
type: debian
install:
container-instructions: |
RUN opam install base stdio utop
ENV OPAM_SWITCH_PREFIX=’/home/opam/.opam/5.0’
ENV CAML_LD_LIBRARY_PATH=’/home/opam/.opam/5.0/lib/stublibs:/home/opam/.opam/5.0/lib/ocaml/stublibs:/home/opam/.opam/5.0/lib/ocaml
ENV OCAML_TOPLEVEL_PATH=’/home/opam/.opam/5.0/lib/toplevel
ENV MANPATH=’:/home/opam/.opam/5.0/man
ENV PATH=’/home/opam/.opam/5.0/bin:/usr/local/sbin:/usr/local/bin:/usr/sbin:/usr/bin:/sbin:/bin
filename: snippet.ml
execute: utop -require base -require stdio snippet.ml
The Fortran configuration:
prompt: |
Use Fortran 90. Some tips:
Always begin each scope with implicit none, pick explicit kinds via
selected_*_kind, and declare proper lengths-character(len=*) is legal only for
dummy arguments, not locals. Strings are blank-padded: call len_trim before
iterating, and store dynamic text in deferred-length allocatables
(character(len=:), allocatable :: s). List-directed read(*,*) arr does not
auto-size arrays; read a count first, then allocate and read, or tokenize a
line manually. When translating 0-based formulas (heaps, bit positions)
remember Fortran arrays default to 1-based; if you want 0-based, declare lower
bounds. Use real literals (2.0d0, 1.0_rk) to avoid silent integer division,
and guard against overflow when exponentiating integers. For frequency tables,
allocate an array or use findloc; Fortran lacks native dicts/sets, so you must
implement search yourself. Prefer array intrinsics (sum, count, pack) over
hand-rolled loops, and keep helper procedures inside a contains section or
module so interfaces are explicit. return inside the main program is
non-idiomatic; use structured blocks or stop. Never print interactive prompts
in batch solutions; just read, compute, and write.
install: apt-get install -y gfortran
filename: snippet.f90
compile: gfortran -o snippet.out snippet.f90
execute: ./snippet.out
Appendix D Training and Results
D.1 Choosing hyperparameters
Before picking the hyperparameters described in § 3.3, we investigated other values by training the Qwen 3 4B model on a previous version of Ag-Codeforces-X. We trained two models for each of Lua, Julia and R, using a linear learning rate schedule with the same learning rate. We decided against it since some of the runs degraded the model’s capabilities, unlike any of the runs we did with a cosine decay schedule. We also compared between GRPO group sizes of 16, 32 and 64 by training the same model on Lua. We observed that in some runs with group size 16, the model improved significantly less than at higher group sizes. Since we found no significant difference between group sizes 32 and 64, we chose the smaller one due to limited resources.
D.2 Training dynamics
Figure 8: Training Qwen3-4B-CF-X, GRPO group batch pass@1.
Figure 9: Training Qwen3-4B-CF-X, test split pass@1.
Figure 10: Training Qwen3-4B-CF-Lua, GRPO group batch and test split pass@1.
Figure 11: Training Qwen3-4B-CF-Julia, GRPO group batch and test split pass@1.
Figure 12: Training Qwen3-4B-CF-R, GRPO group batch and test split pass@1.
Figure 13: Training Qwen3-4B-CF-OCaml, GRPO group batch and test split pass@1.
Figure 14: Training Qwen3-4B-CF-Fortran, GRPO group batch and test split pass@1.
In this section we discuss the measurements we took while training the models. Figure 8 shows the GRPO group batch pass@1 while training the Qwen3-4B-CF-X models. The scores of all the models are broadly correlated with one another, in part due to training them on the same permutation of the training data. Figures 10, 11, 12, 13 and 14 compare the GRPO group pass@1 scores with the pass@1 scores on the test split. We see that the scores on the test split is broadly correlated with the train split rewards. In most cases, we see that the train scores keep increasing until the end of the epoch, together with the test split pass@1 scores, indicating that the model keeps improving until the end of the dataset.
D.3 Hardware and Software Used
We used three machines while working on this paper: B, R1 and R2. R2 was only used to generate completions of trained models for evaluation, while B and R1 were used to train models. Upon publication of the paper, we will publicly release Wandb records of our training runs, which include the duration and the machine used.
B has 2 Intel Xeon Gold 6342 CPUs @ 2.80GHz, 1008 GB of RAM, 4 NVIDIA H100 80GB, and uses Ubuntu 22.04.5 LTS.
R1 has 2 AMD EPYC 9454 48-Core CPUs, 8 NVIDIA H100 80GB (with NVLink connections), 2268 GB of RAM, and uses Ubuntu 22.04.5 LTS.
R2 has 2 Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz, 10 NVIDIA RTX A600, 504 GB of RAM, and uses Ubuntu 22.04.5 LTS.
When developing the Agnostics framework, we used the following major Python libraries: ray v2.46.0, torch v2.6.0, transformers v4.54.1, vllm v0.8.5.post1, datasets v3.4.1, wandb 0.19.11.
Appendix E Bug Taxonomy
E.1 Prompt for Generating Taxonomy
We used the following instructions to generate the bug taxonomy, followed by a list of faulty R programs.
Input: The attached file contains multiple failed R programs (Version 4) with their:
  • Source code
  • Expected output
  • Actual standard output
  • Error messages (where applicable)
Objective: Analyze these program failures systematically to create a comprehensive taxonomy of 10-12 bug themes that categorize the underlying causes of failure.
Instructions:
  1. 1.
    Initial Analysis
    • Read through ALL program examples carefully
    • For each failure, identify the root cause (not just the symptom)
    • Note any patterns or commonalities across failures
  2. 2.
    Taxonomy Development
    • Create 10-12 distinct bug themes that collectively cover all observed failures
    • Each theme should represent a fundamental type of programming error or misconception
    • Themes should be mutually exclusive when possible, but comprehensive in coverage
    • Order themes from most to least frequent (or by logical grouping)
  3. 3.
    For Each Bug Theme, Provide:
    • Theme Name: A concise, descriptive title
    • Description: 2-3 sentences explaining the nature of this bug type
    • Common Symptoms: How these bugs typically manifest (error messages, incorrect output, etc.)
    • Root Causes: The underlying programming mistakes or misconceptions
    • Examples: Reference 2-3 specific programs from the file that exhibit this theme
    • Prevention Tips: Brief advice on how to avoid this type of bug
  4. 4.
    Constraints:
    • Focus on R-specific issues as well as general programming errors
    • Base your taxonomy ONLY on the provided examples
    • You may search online ONLY to understand specific R error messages or function behavior, not for existing bug taxonomies
    • Ensure every failed program in the file can be classified under at least one theme
  5. 5.
    Deliverable Format: Present your taxonomy as a numbered list with clear formatting and comprehensive coverage of all observed failure patterns. Supply a short explanation for each theme in your taxonomy.
The prompt produced a taxonomy of 11 bug categories. We edited these categories and selected 7 categories relevant to us, shown in §E.2.
E.2 Bug Taxonomy Used For Analysis
The following categories represent the prevalent themes of programming errors we use in our analysis of bugs in model-generated code. They cover the full spectrum of parse, runtime, and logical failures typically encountered in programming. The themes are not mutually exclusive; we allow a program to have more than one themes.
  1. 1.
    Syntax and Typographical Errors: Missing commas, mismatched parentheses, or other typos that cause compile-time parse errors.
  2. 2.
    Input Reading and Parsing Errors: Mis-reading or mis-parsing input, leading to empty or malformed variables and subsequent failures.
  3. 3.
    Uninitialized Variables: References to variables never defined, causing undefined behavior or runtime faults.
  4. 4.
    Data Type and Conversion Errors: Incorrect casting or type misuse that triggers type errors, warnings, or incorrect results.
  5. 5.
    Function Misuse and Missing Libraries: Invocations of non‑existent or mis‑parameterized functions, or missing imports/libraries, causing errors.
  6. 6.
    Algorithmic Logic Flaws: Programs that compile and run but produce wrong answers due to faulty logic or conditions.
  7. 7.
    Output Formatting and Presentation Errors: Correct computational results, but incorrect due to formatting issues (e.g. missing newlines/spaces or output spec violations).
E.3 Radar Charts for All Programming Languages
Figures 15, 16, 17, 18, 19 show the error theme charts for all the programming languages we trained a model on. Figure 18 is the same as Figure 6 from the main body; we repeated it here for convenience.
Figure 15: Radar chart of Lua error themes for Qwen3-4B and Qwen3-4B-CF-Lua.
Figure 16: Radar chart of Julia error themes for Qwen3-4B and Qwen3-4B-CF-Julia.
Figure 17: Radar chart of R error themes for Qwen3-4B and Qwen3-4B-CF-R.
Figure 18: Radar chart of OCaml error themes for Qwen3-4B and Qwen3-4B-CF-OCaml.
Figure 19: Radar chart of Fortran error themes for Qwen3-4B and Qwen3-4B-CF-Fortran.
Appendix F Reproducibility Checklist
This paper:
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced: yes
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results: yes
  • Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper: yes
Does this paper make theoretical contributions? No.
Does this paper rely on one or more datasets? Yes.
  • A motivation is given for why the experiments are conducted on the selected datasets: yes.
  • All novel datasets introduced in this paper are included in a data appendix: no.
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: yes.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations: yes.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available: yes.
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisfying: N/A.
Does this paper include computational experiments? Yes.
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting: yes.
  • Any code required for pre-processing data is included in the appendix: no.
  • All source code required for conducting and analyzing the experiments is included in a code appendix: no.
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: yes.
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from: no.
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results: yes.
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks: yes.
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics: yes.
  • This paper states the number of algorithm runs used to compute each reported result: yes.
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information: no.
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank): no.
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments: yes.

### v2

@@
-Agnostics: Learning to Code in Any Programming Language via Reinforcement with a Universal Learning Environment
+Agnostics: Learning to Synthesize Code in Any Programming Language with a Universal Reinforcement Learning Environment
 Abstract
-Large language models (LLMs) already excel at writing code in high-resource languages such as Python and JavaScript, yet stumble on low-resource languages that remain essential to science and engineering.
+Large language models (LLMs) excel at writing code in high-resource languages such as Python and JavaScript, yet stumble on low-resource languages that remain essential to science and engineering.
 Besides the obvious shortage of pre-training data, post-training itself is a bottleneck: every new language seems to require new datasets, test harnesses, and reinforcement-learning (RL) infrastructure.
@@
 Concretely, we (i) use an LLM to rewrite existing unit-test datasets into an I/O format, (ii) supply a short configuration that tells the verifier how to compile and run a target language, and (iii) apply reinforcement learning with verifiable rewards (RLVR) in a robust code execution environment.
-Applied to five low-resource languages—Lua, Julia, R, OCaml, and Fortran—Agnostics (1) improves Qwen-3 4B to performance that rivals other 16B–70B open-weight models; (2) scales cleanly to larger and diverse model families (Qwen-3 8B, DeepSeek Coder 6.7B Instruct, Phi 4 Mini); and (3) for ${\leq}16$B parameter models, sets new state-of-the-art pass@1 results on MultiPL-E and a new multi-language version LiveCodeBench that we introduce.
-We will release the language-agnostic training datasets (Ag-MBPP-X, Ag-Codeforces-X, Ag-LiveCodeBench-X), training code, and ready-to-use configurations, making RL post-training in any programming language as simple as editing a short YAML file.
+Applied to five low-resource languages—Lua, Julia, R, OCaml, and Fortran—Agnostics (1) improves Qwen-3 4B to performance rivaling other 16B–70B open-weight models; (2) scales to larger and diverse model families (Qwen-3 8B, DeepSeek Coder 6.7B Instruct, SmolLM 3, Phi 4 Mini); and (3) for ${\leq}16$B parameter models, sets new state-of-the-art pass@1 results on MultiPL-E and a new multi-language version of LiveCodeBench which we introduce.
+We release the language-agnostic training datasets (Ag-MBPP-X, Ag-Codeforces-X, Ag-LiveCodeBench-X), training code, and ready-to-use configurations, making RL training for any programming language as easy as a few lines of YAML.
 1 Introduction
-Figure 1: Overview: Agnostics Data Preparation and Training.
-(1) We reformulate existing coding datasets to our format.
-(2) We adapt the language-agnostic datasets to a particular programming language.
-(3, 4) We reinforce coding via Group Relative Policy Optimization (Shao et al.
-2024; DeepSeek-AI et al.
-2025), verifying the programs in our code execution sandbox.
-Large language models (LLMs) are remarkably good at programming tasks, especially when codinng in high-resource programming languages such as Python and JavaScript.
-Their proficiency in low-resource programming languages, such as Fortran, Julia, and others, is significantly more limited.
-This gap appears both on benchmarks (Cassano et al.
-2023) and in popular discourse.
-However, many low-resource languages are widely used in computational science (e.g., Fortran), medicine (e.g., Mumps), engineering (e.g., Matlab), and others sectors.
-For programmers in these sectors to be able to take full advantage of LLMs, we need methods to make LLMs better at the languages that they use.
-The capability gap between high-resource and low-resource languages occurs for two reasons.
-The first one is that there is vastly more training data for some programming languages.
-For example, The Stack V2 (Lozhkov et al.
-2024a), the largest public training corpus of code, has ${\approx}200$GB of Python but only ${\approx}2$GB of Julia and Fortran.
+Large language models (LLMs) are remarkably good at programming tasks, especially when coding in high-resource programming languages such as Python and JavaScript.
+Their proficiency in low-resource programming languages, such as Fortran, Julia, and others, is far more limited.
+This gap appears both on benchmarks (Cassano et al., 2023) and in popular discourse.
+Many low-resource languages are adapted to and widely used in particular sectors such as computational science (e.g., Julia, Fortran), medicine (e.g., Mumps), data science (e.g., R), and others.
+Methods for improving LLMs on such languages would help programmers in these sectors truly take advantage of LLMs.
+The capability gap between high-resource and low-resource programming languages occurs for two reasons.
+First, there is vastly more training data for some languages.
+For example, The Stack V2 (Lozhkov et al., 2024a), the largest public training corpus of code, has ${\approx}200$GB of Python but only ${\approx}2$GB of Julia and Fortran.
 Thus pretraining on code makes models significantly better at Python.
-A more subtle reason is the availability of post-training datasets and techniques.
+A subtler reason is the availability of post-training datasets and techniques.
 Contemporary LLMs are developed with an extensive post-training process that relies on (a) high-quality curated data for supervised fine-tuning, and (b) carefully designed environments for reinforcement learning, which must be able to execute and verify model-generated solutions.
-Both of these require significant human expertise, which is hard to find for low-resource languages.
-This paper aims to dramatically facilitate post-training LLMs on low-resource programming languages, working towards closing the resource gap.
-Our key idea is that for a significant class of programming tasks, correctness can be specified as a property not of individual functions or code snippets, but of the entire program’s observable behavior (e.g., I/O).
-Furthermore, if its correctness can be tested with another program, we can use such a verifier together with appropriate problems and test cases to make a universal reinforcement learning environment which can be adapted to any programming language.
-In fact, the language used to implement the verifier and the one being learned are entirely independent.
-This approach matches how some existing post-training datasets are formulated (even though they are intended for Python/C++).
-We can also use LLMs to reformulate language-specific datasets into this format.
-Based on this insight, our approach, Agnostics, works as follows (Figure 1).
-1) We use an LLM to reformulate language-specific datasets into our standard language-agnostic format.
-2) We generate prompts and a configured verifier targeting a particular language, with a small (4-5 line) configuration.
+Both of these require significant human expertise, which is hard to find for low-resource programming languages.
+Our goal in this work is to facilitate post-training LLMs on low-resource programming languages, working towards closing the resource gap.
+Our key idea is that for a large class of programming tasks, correctness can be stated as a property not of functions or code snippets, but of the entire program’s observable behavior (e.g., I/O).
+Furthermore, if its correctness can be tested with a verifier program, such a verifier with appropriate problems and test cases can be used to make a universal reinforcement learning environment which can be instantiated for nearly any programming language.
+In fact, the verifier’s implementation language is independent from the one being learned.
+This approach matches the formulation of some existing post-training datasets (even if they are intended for Python/C++), and we can reformulate language-specific datasets into this format with LLMs.
+Our approach, Agnostics, works based on this insight as follows (Figure 1).
+1) We use an LLM to reformulate language-specific datasets into our uniform language-agnostic format.
+2) To target a particular language, we generate prompts and instantiate the verifier based on a small (4-5 line) configuration file.
 3) We apply reinforcement learning with verified rewards (RLVR) using a robust, language-agnostic execution sandbox that we develop.
@@
 Contributions
-We make the following contributions.
   1. 1.
-    Agnostics, a post-training pipeline for coding in arbitrary programming languages, including low-resource ones;
+    Agnostics, a post-training pipeline for coding in arbitrary programming languages;
   2. 2.
-    The best-performing open-weights models for Lua, R, Julia, OCaml and Fortran at the ${\leq}16$B parameter scale;
+    The best-performing open-weights ${\leq}16$B models for Lua, R, Julia, OCaml and Fortran;
   3. 3.
-    3 Agnostics datasets Ag-MBPP-X, Ag-Codeforces-X, and Ag-LiveCodeBench-X, based on MBPP (Austin et al.
-2021), Open-R1 Codeforces (Penedo et al.
-2025) and LiveCodeBench (Jain et al.
-2024a) respectively.
+    Three Agnostics datasets: Ag-MBPP-X, Ag-Codeforces-X, and Ag-LiveCodeBench-X, based on MBPP (Austin et al., 2021), Open-R1 Codeforces (Penedo et al., 2025) and LiveCodeBench (Jain et al., 2024b) respectively.
   4. 4.
-    A small and highly-optimized Agnostics training framework, implementing GRPO and including a parallel code execution sandbox, sampling, rewards computation, and model back-propagation.
-Artifacts and code necessary to verify the results we present are listed at agnostics.abgru.me. The training framework will be publicly released together with the formal publication of this work.
+    A small and carefully designed Agnostics training framework, including a parallel code execution sandbox, sampling, rewards computation, GRPO, and model back-propagation.
+Our leaderboard, data, code, and artifacts are referenced at agnostics.abgru.me and in Reproducibility Statement.
+Figure 1: Overview: Agnostics Data Preparation and Training.
+(1) We reformulate existing coding datasets to our format.
+(2) We adapt the language-agnostic datasets to a particular programming language.
+(3, 4) We reinforce coding via Group Relative Policy Optimization (Shao et al., 2024; DeepSeek-AI et al., 2025), verifying the programs in our code execution sandbox.
 2 Background and Related Work
+Data Scarcity for Low-Resource Languages
 […]
-Figure 2: High-resource languages (left) dominate the Stack v2. We are interested in low-resource languages (right).
-Data Scarcity for Low-Resource Languages
-LLM have been pretraining on code for several years, both because LLMs are widely used for practical programming tasks, and because pretraining on code improves their general reasoning abilities (Ma et al.
-2023).
-In addition to general-purpose LLMs, there are also code-specialized models that are either trained exclusively on code (e.g., Xu et al.
+Figure 2: High-resource (left) and low-resource (right) languages in the Stack v2.
+General-purpose LLMs have been pretrained on code for several years, both because LLMs are widely used for practical programming tasks, and because pretraining on code improves their general reasoning abilities (Ma et al., 2023).
+Code-specialized models may be trained from a general-purpose model checkpoint or exclusively on code (e.g., Xu et al.
 (2022); Allal et al.
-(2023)) or trained on code starting from a general-purpose model checkpoint (e.g., Rozière et al.
 (2023)).
 However, the publicly available pretraining data for code is heavily skewed toward a handful of programming languages.
-For example, consider The Stack V2 (Lozhkov et al.
-2024b), which is the largest public code pretraining dataset, with code from GitHub and dozens of other sources.
-The Stack V2 dominated by a handful of programming languages: just 10 of 619 languages account for over 90% of the dataset.
-We are interested in developing models for low-resource languages that each account for less than 0.5% of the publicly available code (figure 2).
+E.g., consider The Stack V2 (Lozhkov et al., 2024b), the largest public code pretraining dataset, with code from GitHub and dozens of other sources.
+The Stack V2 is dominated by relatively few programming languages: just 10 of 619 languages account for over 90% of the dataset.
+We want to develop models for low-resource programming languages that each account for ${\leq}0.5\%$ of publicly available code (figure 2).
 We can imagine working around this data scarcity in a few ways.
-However, previous work shows that up-sampling low-resource languages during pretraining only leads to small benchmark improvements (Orlanski et al.
-2023), while fine-tuning on the pre-training data for low-resource languages has negligible impact (Cassano et al.
-2024).
+However, previous work shows that up-sampling low-resource languages during pretraining only leads to small benchmark improvements (Orlanski et al., 2023), while fine-tuning on the pre-training data for low-resource languages has negligible impact (Cassano et al., 2024).
 Thus, it is not clear how to make further gains from existing natural data.
 Synthetic Data for Low-Resource Languages
-If natural data is not available for a task, it is possible to use LLMs to generate synthetic fine-tuning data (Wang et al.
-2023), and there are many techniques for building code-centric supervised fine-tuning datasets (e.g.
+If natural data is not available for a task, it is possible to use LLMs to generate synthetic fine-tuning data (Wang et al., 2023), and there are many techniques for building code-centric supervised fine-tuning datasets (e.g.
 Luo et al.
 (2023); Wei et al.
-(2024c, b)) that work remarkably for Python.
+(2024c; b)) that work remarkably for Python.
 Although these approaches could in principle be applied to any programming language, Cassano et al.
@@
 (2024a)), synthetic tasks and code for low-resource programming languages are low-quality, and models fine-tuned on them perform poorly.
-MultiPL-T (Cassano et al.
-2024), similar to TransCoder-ST (Roziere et al.
-2021) and CMTrans (Xie et al.
-2024), couples synthetic data generation with verification using rejection sampling: it generates up $n$ candidate programs in a target, low-resource language and only fine-tune models on generations that pass hidden unit tests that it translates from Python.
+MultiPL-T (Cassano et al., 2024), similar to TransCoder-ST (Roziere et al., 2021) and CMTrans (Xie et al., 2024), couples synthetic data generation with verification using rejection sampling: it generates up $n$ candidate programs in a target, low-resource language and only fine-tune models on generations that pass hidden unit tests that it translates from Python.
 However, the MultiPL-T approach has two significant limitations.
@@
 In MultiPL-T, ${\approx}30\%$ of prompts produce a working program for $n\in\{50,100\}$ attempts, and the rest are discarded.
-We train on much harder problems, and estimate that rejection sampling would require an order of magnitude more resources to achieve a comparable acceptance rate (§ A).
-(2) For each low-resource language of interest, MultiPL-T requires the user to write a little compiler to translate test cases and function signatures from Python to the target language.
-MultiPL-T only supports a limited set of built-in Python types (e.g., no classes) and it dictates that all Python types and values must faithfully map to the target language.
-However, depending on the problem and language, the natural way to represent data may not map cleanly to Python types.
+We train on much harder problems, and estimate that rejection sampling would require an order of magnitude more resources for a comparable acceptance rate (§ 4.3).
+(2) For each low-resource language of interest, MultiPL-T requires writing a little compiler to translate test cases and function signatures from Python to the target language.
+MultiPL-T only supports a limited set of built-in Python types (e.g., no classes) and dictates that all Python types and values must faithfully map to the target language.
+However, depending on the problem and language, the natural data representation may not map cleanly to Python.
 This can lead to peculiar, unidiomatic translations that require deep language expertise to get right.
-The Agnostics approach is significantly easier to use than MultiPL-T, and only requires the user to know how to compile and run a program in the target language from the shell.
+The Agnostics approach is far easier to use than MultiPL-T, and only requires the user to know how to compile and run a program in the target language from the shell.
 Reinforcement Learning on Coding Tasks
-DeepSeek R1 (DeepSeek-AI et al.
-2025) popularized RL on LLMs with rule-based rewards, instead of learned reward models.
-R1 reports applying RL to coding tasks but does not disclose further dataset details.
-A number of papers apply RL to the NL to code task (Zeng et al.
-2025; Gehring et al.
-2024; Jain et al.
-2025).
-These techniques, which target Python, show that reinforcement learning can improve LLM capabilities beyond what is possible with just supervised fine-tuning.
-However, the key benefit of reinforcement learning is that it can train a model to do tasks for which high-quality data for supervised fine-tuning is unavailable.
-There are recent examples of using RL for code optimization (Du et al.
-2025; Nichols et al.
-2024), resolve GitHub, issue resolution (Wei et al.
-2025), and iterative development (Zhou et al.
-2025).
-These papers target tasks in high-resource languages—C++, Java, and Python—whereas Agnostics targets several low-resource languages.
+DeepSeek R1 (DeepSeek-AI et al., 2025) popularized RL on LLMs with rule-based rewards, instead of learned reward models.
+R1 reports applying RL to coding tasks without further dataset details.
+A number of papers apply RL to the NL to code task (Zeng et al., 2025; Gehring et al., 2024; Jain et al., 2025).
+These techniques target Python and show that RL can improve LLM capabilities beyond what supervised fine-tuning allows alone.
+However, the key benefit of RL is that it can train a model to do tasks for which high-quality supervised fine-tuning data is unavailable.
+There are recent examples of using RL for code optimization (Du et al., 2025; Nichols et al., 2024), resolve GitHub, issue resolution (Wei et al., 2025), and iterative development (Zhou et al., 2025).
+These papers target tasks in high-resource languages (C++, Java, and Python) whereas Agnostics targets several low-resource languages.
 3 The Agnostics Approach
-Our approach comprises (1) a data preparation stage which reformulates language-specific programming tasks to be language-agnostic, and retargets language-agnostic datasets to a programming language of interest (1, 2 in figure 1); and (2) the training stage which uses the GRPO algorithm with an efficient, language-agnostic verification framework (3, 4 in figure 1).
-Our central idea is to have tasks ask for programs which behave in a particular way.
+Our approach comprises (1) a data preparation stage which reformulates language-specific programming tasks to be language-agnostic, and retargets language-agnostic datasets to a programming language of interest (1, 2 in figure 1); and (2) the training stage which uses the GRPO algorithm and an efficient, language-agnostic verification framework (3, 4 in figure 1).
+Our tasks ask for programs by describing their behavior.
 The test cases are samples of this behavior, and a verifier program can check if a solution behaves according to the sample.
 In this paper, we limited ourselves to working with tasks asking for programs which read data from the standard input, compute a unique answer, and write it to the standard output.
-Hence, the datasets we prepared share the same verifier.
-# Write a python function to identify non-
-# prime numbers.
+Hence, the datasets we prepared share one verifier.
+# Write a python function to
+# identify non-prime numbers.
 def is_not_prime(n):
@@
 assert is_not_prime(10) == True
-(a) A task prompt and associated tests (in gray) from MBPP.
-Instruction: Given an integer NN ($N\geq 2$), determine whether it is a non-prime number.
+(a) An MBPP task prompt and associated tests (in gray).
+Instruction: Given an integer NN ($N\geq 2$), determine if it is a non-prime number.
 Output ‘True’ if the number is non-prime, ‘False’ otherwise.
@@
 3.1 Dataset Preparation
-Some datasets, like Open-R1 Codeforces (Penedo et al.
-2025), already define tasks in the desired I/O style.
+Some datasets, like Open-R1 Codeforces (Penedo et al., 2025), already define tasks in the desired I/O style.
 More commonly, however, code datasets provide a set of unit tests.
@@
 These datasets can be easily translated into the I/O format.
-To make such problems language‑agnostic and compatible with our verifier, we prompt an LLM to reformulate each task so that the program communicates exclusively via plain‑text standard in and standard out.
+To make such problems language-agnostic and compatible with our verifier, we prompt an LLM to reformulate each task so that the program communicates exclusively via plain‑text standard in and standard out.
 We ask the model to spell out concrete I/O conventions—number of decimal places, newline versus comma separators, ordering of values, and so on—so that the expected behavior is unambiguous.
 Figure 3(b) shows the reformulated example.
-§ B has the instruction we use to reformulate MBPP; other datasets might require small changes to the prompt.
+§ A presents the prompt we use to reformulate MBPP; other datasets might require small changes to the prompt.
 3.2 Programming Language Preparation
-install: apt-get install -y r-cran-tidyverse
-filename: snippet.R
-execute: Rscript snippet.R
-prompt: |
-Use R version 4.
-Use readLines(con = file("stdin"))‘ to
-read input from stdin. Optionally, use
-the n argument to read the first n
-lines. For example:
-‘‘‘r
-input <- readLines(con = file("stdin"), n = 1)
-n <- as.integer(input)
-print(n) # print the first line of input
-Also, use cat to print output to stdout.
-For example:
-‘‘‘r
-cat(n)
-Please do not use print to print output.
-Figure 4: An Agnostics configuration snippet for R.
-To prepare a new language, we author a small configuration file that serves two purposes.
+To prepare a new language, we author a small configuration file with two purposes.
 First, it defines a prompt prefix (prepended to each problem by the trainer) which instructs the model to produce code in the target language.
 Second, the configuration file specifies the shell commands to install the language toolchain and run code.
-Figure 4 shows the configuration for the R language, with an atypically large prompt due to quirks of its I/O APIs.111 There are three ways to run R, three I/O APIs, and only one portable way to read from standard in.
-A good prefix should help avoid the most common errors, but determining which errors are common can be time-consuming.
-In our experience, if the model starts from ${\geq}5\%$ accuracy, then the prefix only needs to say that the solution must be in language $L$.
-(This varies with GRPO group size, see § 3.3.) On the other hand, when starting from near-zero accuracy, a prefix like in figure 4 can help prevent common mistakes.
-For OCaml and Fortran, we asked a capable LLM (OpenAI o3) for advice based on several faulty generations.
+In our experience, a prompt prefix simply asking for a solution in language $L$ is enough for more widespread languages with ${\geq}5\%$ base accuracy.
+However, when starting from near-zero accuracy, a longer prefix can help prevent common mistakes.
+E.g., our R language configuration (figure 4) features a longer prompt explaining the quirks of I/O APIs in R.
+[^1] There are 3 ways to run R, 3 I/O APIs, and only one portable way to read from standard in.
+If a model barely knows a programming language, a good prefix can help it.
+Still, writing the prefix takes manual effort.
+For OCaml and Fortran, we let a base model generate several faulty snippets, and asked a capable LLM (OpenAI o3) for advice based on the snippets with the following prompt.
 What follows are several Fortran programs.
@@
 Distill this advice into 10-20 sentences.
-We used the resulting instructions verbatim (§ C), but only for training and not to evaluate models on OCaml or Fortran.
+We use the resulting instructions verbatim (see § B) when training models.
+The prefix only needed to slightly raise the model’s train split performance; base accuracy as low as $0.09\%$ was enough for the model to start learning (see § 4.3).
+Configuring the two languages took 1 hour each.
+install: apt-get install -y r-cran-tidyverse
+filename: snippet.R
+execute: Rscript snippet.R
+prompt: |
+Use R version 4. Use readLines(con = file("stdin"))‘ to read from
+stdin. Use the n argument to read the first n lines. For example:
+‘‘‘r
+input <- readLines(con = file("stdin"), n = 1)
+n <- as.integer(input)
+cat(n) # print the first line of input
+Also, please remember to use cat to print output.
+Figure 4: An Agnostics configuration snippet for R (slightly rephrased for presentation).
 3.3 Trainer and Code Execution
-The Agnostics trainer uses the Group-Relative Policy Optimization (GRPO) reinforcement learning algorithm (Shao et al.
-2024), with verifiable rewards (DeepSeek-AI et al.
-2025), and some common tweaks to further improve its efficiency (Yu et al.
-2025).
-We couple the algorithm with a language-agnostic code execution framework that is designed to be robust and efficient.
+The Agnostics trainer uses the Group-Relative Policy Optimization (GRPO) reinforcement learning algorithm (Shao et al., 2024), with verifiable rewards (DeepSeek-AI et al., 2025), and further common tweaks to improve its efficiency (Yu et al., 2025).
+We couple the algorithm with a language-agnostic code execution framework designed to be robust and efficient.
 Trainer
@@
 Let $(x,\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K})\sim\mathcal{D}$ be a dataset of language-agnostics tasks, where $x$ is the task prompt and $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$ is the set of I/O examples.
-Let $L$ be a language configuration $L$ (e.g., figure 4).
-From the behavior policy $\pi_{\theta_{\text{old}}}$, we sample a group $G$ of candidate responses
-$\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(y_{i}\mid L.\mathsf{prompt},x).$
-We assign each candidate a reward $R_{i}$, with $R_{i}=1$ if the execution environment (described later) verifies that the extracted program’s behavior is consistent with the I/O examples $(\mathit{in}_{k},\mathit{out}_{k})$, and $R_{i}=0$ otherwise.
-We turn group rewards into sequence-level advantages
-$\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}(\{R_{j}\}_{j=1}^{G})}{\operatorname{std}(\{R_{j}\}_{j=1}^{G})}.$
-We then update the policy with the objective
-$\mathcal{L}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Bigl{(}r_{i,t}(\theta)\hat{A}_{i},c_{i,t}(\theta)\hat{A}_{i}\Bigr{)},$
-where
-$c_{i,t}(\theta)$ $=\text{clip}\bigl{(}r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\bigr{)},$
-$r_{i,t}(\theta)$ $=\frac{\pi_{\theta}(y_{i,t}\mid L.\mathsf{prompt},x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid L.\mathsf{prompt},x,y_{i,<t})}.$
-We omit the KL-divergence term, similar to Yu et al. (2025).
-We considered and decided against a reward for a partially-correct answer.
-We tried to reward the model for code which runs without errors but produces wrong output or for code which only passes the public tests.
+Let $P$ be $L.\texttt{prompt}$ from a language configuration $L$ (e.g., figure 4).
+From the behavior policy $\pi_{\theta_{\text{old}}}$ we sample a group $G$ of candidate responses $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}({\cdot}\mid{P,x}).$ We assign each candidate a reward $R_{i}$, with $R_{i}=1$ if the execution environment (described later) verifies that the extracted program behaves as in the I/O examples $(\mathit{in}_{k},\mathit{out}_{k})$ and $R_{i}=0$ otherwise.
+We turn group rewards into sequence-level advantages $\hat{A}_{i}$, and update the policy with the objective
+$\mathcal{L}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{\{(x,\_)\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid P,x)\}}$
+$\quad\Biggl[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Bigl(\text{clip}\bigl(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\bigr)\hat{A}_{i},r_{i,t}(\theta)\hat{A}_{i}\Bigr)\Biggr],$
+$\text{where}\qquad r_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid P,x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid P,x,y_{i,<t})},\qquad\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}(\{R_{j}\}_{j=1}^{G})}{\operatorname{std}(\{R_{j}\}_{j=1}^{G})}.$
+We omit the KL-divergence term, similar to Yu et al.
+(2025).
+We also considered and decided against a reward for a partially-correct answer (see § C.3).
+We tried to reward the model for code which runs without errors but produces wrong output or for code which only passes the public tests (if there are any).
 In both cases the models were very likely to learn how to exploit the reward, e.g., by producing empty programs or by hard-coding the public tests (and claiming to produce a “draft answer”).
 Code Execution
-Our verifier is a language-agnostic code execution sandbox which (1) extracts a program from each candidate; (2) compiles it if needed; and (3) tests it on the I/O examples $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$.
+Our verifier, a language-agnostic code execution sandbox, (1) extracts a program from each candidate; (2) compiles it if needed; and (3) tests it on the I/O examples $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$.
 To extract the code, we instruct the model to put it in a Markdown block, which all major instruction-tuned models do by default.
 Since we rely on the native format of the model, we do not need to train the model with a format reward.
-This guarantees that the increases in rewards that we see during training are genuine improvements and not merely the result of the model learning to format correctly.
-For each language, we build and cache an OCI container using the configuration $L$.
+This guarantees that the increases in rewards we see during training are real improvements and not merely the result of the model learning to format correctly.
+For each language, we build and cache an OCI (2025) container using the configuration $L$.
 To build the container, we install the language compiler and runtime (the script $L.\texttt{install}$), and include a generic execution harness which runs and tests candidate programs.
-The execution harness runs continuously in the container, waiting to receive a triple with the candidate program, the set of input/output examples, and timeouts.
+The execution harness runs continuously in the container, waiting for triples with the candidate program, the set of input/output examples, and timeouts.
 The harness (1) writes the program to disk (to $L.\texttt{filename}$), (2) compiles it if needed ($L.\texttt{compile}$), (3) runs it on each received input ($L.\texttt{execute}$), and verified that it produces the expected output.
@@
 Finally, to improve compile times, our execution environment mounts a RAM disk in each container.
-Compilation may be slow due to creating many intermediate files, and indeed some large C++ projects, e.g., Firefox, recommend using a RAM disk to speed up their builds (Firefox 2025).
+Compilation may be slow due to creating many intermediate files, and indeed some large C++ projects, e.g., Firefox, recommend using a RAM disk to speed up their builds (Firefox, 2025).
 Implementation
-We implement the trainer and execution environment with Ray (Moritz et al.
-2018), which facilitates multiprocessing and distributed computing.
-In particular, the framework allows distributing the training over a network of heterogeneous nodes, which allows running the trainer on a node specialized for GPU work and the execution environment on a node specialized for CPU work.
-Ray also lets us easily have two separate inter-communicating processes, one for group generation and another for verification and loss computation, simply by defining two actor classes.
-Running the two concurrently significantly speeds up training, as we found that they take a roughly comparable amount of time.
-The execution environment is also implemented as an actor, with one Python asyncio task per container.
-Using tasks instead of actors helps lower how much inter-process communication and data copying is needed.
+We implement the trainer and execution environment with Ray (Moritz et al., 2018), which facilitates multiprocessing and distributed computing.
+In particular, Ray lets us distribute the training over a network of heterogeneous nodes, which allows running the trainer on a node specialized for GPU work and the execution environment on a node specialized for CPU work.
+Ray also lets us easily separate group generation and loss computation into inter-communicating processes.
+Running the two in parallel significantly speeds up training, as we found that they take a roughly comparable amount of time.
+The execution environment is also an actor and manages containers with Python asyncio coroutines, not actors, to minimize inter-process data copying.
 Hyperparameters
-For our training, we use the AdamW optimizer (Loshchilov and Hutter 2019) with a learning rate of $5\times 10^{-6}$ and a cosine decay schedule with a warmup of 0.1 epochs.
+We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $5\times 10^{-6}$ and a cosine decay schedule with a warmup of 0.1 epochs.
 We process 4 prompts in each batch, with group size 32 per prompt.
-We generate with temperature 0.7.
-We train hybrid reasoning models with reasoning disabled.
-(We still often observe generations with reasoning-like text, either before the answer or in comments.)
+When training we use temperature 0.7 and disable reasoning in hybrid models.
+(Many generations still show reasoning-like text, either before the answer or in comments.) We describe how we chose the hyperparameter values and how we evaluted different values in § C.1.
 4 Evaluation
-To evaluate our approach and codebase, we train and benchmark multiple models on 5 low-resource programming languages.
-We measure pass@1 accuracy with reasoning disabled and 20 samples per prompt at temperature 0.2.
+To evaluate Agnostics, we train and benchmark models on 5 low-resource programming languages.
+We measure pass@1 accuracy with reasoning disabled, 20 samples per prompt at temperature 0.2.
+We trained the models for 1 epoch, unless specified otherwise.
 4.1 Training datasets
-Ag-Codeforces-X, the main dataset we use for training, was created based on competitive programming problems from the Open-R1 Codeforces dataset (Penedo et al.
-2025).
+Ag-Codeforces-X, the main dataset we use for training, was created based on competitive programming problems from the Open-R1 Codeforces dataset (Penedo et al., 2025).
 Few adjustments were necessary, since the problems already specified programs and tests using standard I/O.
 The train split contains 5369 problems.
-See § B for more details.
+Ag-MBPP-X, the other training dataset we use, was created from MBPP as explained in § 3.1.
+Both datasets were analyzed to ensure no contamination with benchmarks we use. See § A for a more detailed description of how the datasets were prepared.
 4.2 Benchmarks
 We evaluate Agnostics with the following benchmarks.
-MultiPL-E (Cassano et al.
-2023) is a well-established benchmark which has been frequently used to evaluate the performance of new LLMs on a broad set of languages (e.g., Kimi Team (2025); Yang et al.
+MultiPL-E (Cassano et al., 2023) is a well-established benchmark, frequently used to evaluate the performance of new LLMs on a broad set of languages (e.g., Kimi Team (2025); Yang et al.
 (2025); Grattafiori et al.
@@
 (2025)).
-MultiPL-E was prepared by compiling HumanEval (Chen et al.
-2021) prompts and unit tests from Python to each target language.
+MultiPL-E was prepared by compiling HumanEval (Chen et al., 2021) prompts and unit tests from Python to each target language.
 Each MultiPL-E programming language requires a ${\approx}500$ LOC prompt and test translator, considerably more effort than writing an Agnostics configuration file.
 A major limitation of MultiPL-E is being too easy for frontier models.
-With Python, frontier models are now evaluated on solving programming contest problems (Jain et al.
-2024a); no multi-language benchmarks are as challenging.
+With Python, frontier models are now evaluated on solving programming contest problems (Jain et al., 2024b); no multi-language benchmarks are as challenging.
 Ag-LiveCodeBench-X, a contribution of this paper, is a new multi-language benchmark derived from LiveCodeBench.
-LiveCodeBench 5.0 has 880 problems, of which 381 have Python starter code and test cases.
-The remaining 499 problems do not use starter code and instead use standard I/O to specify and test solutions.
-Hence we used these problems to transform LiveCodeBench into an Agnostics dataset.
-Accordingly, benchmarking a new programming language with Ag-LiveCodeBench-X is straightforward: we can reuse the language configurations and execution environment from our trainer (§ 3.3).
-Importantly, the problems from LiveCodeBench and thus also Ag-LiveCodeBench-X are filtered out from our training sets.
-Moreover, our results show that Ag-LiveCodeBench-X is significantly harder than MultiPL-E.
+(The benchmark has no overlap with our training data; see § A.) LiveCodeBench 5.0 has 880 problems; 381 problems have Python starter code and test cases, while the remaining 499 problems instead use standard I/O to specify and test solutions.
+We used the latter problems to transform LiveCodeBench into an Agnostics dataset.
+Accordingly, benchmarking a new programming language with Ag-LiveCodeBench-X is straightforward: we can reuse our trainer’s language configurations and execution environment (§ 3.3).
+Our results and leaderboard (see § 1) show that Ag-LiveCodeBench-X is far harder than MultiPL-E and is challenging for frontier models as of February 2026.
 4.3 Results
+Model Ag-LCB-X
+X= Lua Julia R
+Llama 3.3 70B Ins 25 22 13
+Qwen 3 32B 22 26 17
+DSC v2 Lite Ins 16B 13 12 9
+[…]
+Qwen 3 8B 11 9 9
+[…]
+[…]
+[…]
+Table 1: Ag-LCB-X pass@1.
 We now present our results.
@@
 Ag-LCB-X stands for Ag-LiveCodeBench-X; we clarify abbreviated model names in the text.
-One table row may present scores of many models; the column for language X shows the score of the model trained on X.
-Table 1: Pass@1 rates on AG-LiveCodeBench-X for Qwen 3 4B models trained on five low-resource languages using Ag-Codeforces-X.
-Each model achieves a SOTA result for an open-weight model with ${\leq}16$B parameters.
-Model Ag-LiveCodeBench-X
-X= Lua Julia R OCaml Fortran
-Llama-3.3-70B-Ins 25 22 13 7 3
-Qwen3-32B 22 26 17 2 1
-DSCv2-Lite-Ins-16B 13 12 9 7 6
+Highlighted rows present our models; note that each cell in such a row presents the score of a different model trained on programming language X.
+We compute the score as explained in § 4.
+SOTA small LLMs for low-resource PLs
+Using Agnostics, we train Qwen 3 4B and 8B on Ag-Codeforces-X specialized to Fortran, Julia, R, Lua, and OCaml.
+The resulting models are state-of-the-art at coding in low-resource programming languages among open-weight models with ${\leq}16$B parameters.
+Model Ag-LCB-X
+X= OCaml Fortran
+Sonnet 4 6 6
+Llama 3.3 70B Ins 7 3
+Qwen 3 32B 2 1
+DSC v2 Lite Ins 16B 7 6
+[…]
+Qwen 3 8B 0 0
 […]
 […]
-Table 2: Pass@1 rates on MultiPL-E for Qwen 3 4B models trained using Ag-Codeforces-X. The significant improvement indicates that they did not overfit to the competitive programming format.
+Table 2: Ag-LCB-X pass@1.
+Agnostics training yields significant improvements on Ag-LiveCodeBench-X (tables 1 and 2).
+(i) On every language, our models match or outperform DeepSeek Coder v2 Lite Instruct (16B), and their performance comes close to or exceeds that of Qwen 3 32B and Llama 3.2 70B.
+(ii) The OCaml and Fortran scores improve from near zero to ${\approx}7\%$ and ${\approx}15\%$, outperforming even some frontier models.
+Importantly, these scores show what the models learned from training: during evaluation we omit the prompt prefix used to facilitate learning (§ 3.1).
+(iii) Finally, the pass@1 scores improve by a factor of 1.5–2x over the base model.
+This indicates that Agnostics can improve models beyond usual training on all the publicly available code data, as we can safely assume that the Qwen models were trained on all such data, like the Llama models (Grattafiori et al., 2024).
 Model MultiPL-E
 X= Lua Julia R
-Qwen3-4B 61 51 36
+Qwen 3 4B 61 51 36
+Qwen 3 8B 63 53 44
 […]
-SOTA small LLMs for low-resource PLs
-Using Agnostics, we train Qwen 3 4B on Ag-Codeforces-X specialized to Fortran, Julia, Lua, OCaml, and R.
-The resulting Qwen3-4B-CF-X models are, to the best of our knowledge, state-of-the art low-resource programming language code models with open-weights and ${\leq}16$B parameters.
-Benchmarking them on Ag-LiveCodeBench-X (table 1), we see significant improvements.
-(i) On every language, the models match or outperform DeepSeek Coder v2 Lite Instruct (16B).
-(ii) On every language, our performance comes close to or even exceeds that of Qwen 3 32B and Llama 3.2 70B.
-(iii) Compared to the base model, Qwen 3 4B, pass@1 improves by a factor of 1.5–2x.
-It is safe to assume that the Qwen models, like the Llama models (Grattafiori et al.
-2024), are trained on all the publicly available coding data; hence, Agnostics improves the models beyond what typical training on such data allows.
-(iv) Finally, on OCaml and Fortran, we improve the performance of Qwen 3 4B from near zero to 7% and 15%.
-Notably, although we train the models with a prompt prefix to facilitate learning (§ 3.1), we do not use this prompt prefix for evaluation.
-Thus the pass@1 scores represent what the model learned, and not information provided in context.
-A potential weakness of our approach is that our models are overspecialized to the competitive programming format, and can only generate programs that use standard I/O.
-We also evaluate with the established MultiPL-E benchmark, which does not use this format, and also find significant improvements (table 2).
+[…]
+[…]
+Table 3: MultiPL-E pass@1.
+Models trained with our approach generalize over the competitive programming format: the improvements are not limited to synthesizing programs using standard I/O.
+To demonstrate this, we evaluate them on the established MultiPL-E benchmark.
+It features problems which ask for Python functions operating on usual Python data structures, and we find that our training also significantly improves the models on such problems (table 3).
 [^2] Note that MultiPL-E does not support OCaml and Fortran.
-This indicates that our models have not overfit to the competition programming format.
-Figure 5: Training Qwen3-4B-CF-X, GRPO batch pass@1.
+We also confirmed that our training does not lower performance on other programming languages (§ C.4).
 Figure 5 shows the GRPO batch pass@1 rates seen when training Qwen3-4B-CF-X.
@@
 Nearly all the models slowly keep improving almost until the dataset end.
-We also observed the train and test split rewards to be correlated with each other (§ D.2).
-Table 3: Pass@1 rates for larger Qwen 3 8B models trained on three languages using Ag-Codeforces-X. Agnostics scales to models of this size.
+We also observed the train and test split rewards to be correlated with each other (§ C.2).
+Figure 5: Group batch pass@1, Qwen3-4B-CF-X.
+Agnostics scales to larger models
+To test if the gains from Agnostics training scale with model size, we train the Qwen 3 8B model on Ag-Codeforces-X specialized to Lua, Julia and R and benchmarked it on Ag-LiveCodeBench-X and MultiPL-E (tables 1 and 3).
+The Qwen3-8B-CF-X models show significant gains on both benchmarks, improving over their 4B counterparts.
+We expect Agnostics to scale to even larger models, with appropriate computing resources.
+However, we found that Agnostics training on Ag-Codeforces-X does not improve two smaller models, Qwen 3 1.7B and Llama 3.2 3B Instruct, perhaps due to the problems being too difficult for models of this size.
+Agnostics works with easier problems
+All of the models we discussed so far were trained on Ag-Codeforces-X.
+To show that Agnostics works with other datasets, we also train models for Julia, Lua, and R using the MBPP training set.
+(§ 3.1 describes how we prepare MBPP.) MBPP problems are trivial compared to the Open-R1 Codeforces problems (see figure 3(a)), and we cannot expect models trained on the MBPP problems to be as good as ones we presented before.
+Still, training on MBPP improves Lua and Julia performance (tables 1 and 3).
+The table shows a small drop in R performance on Ag-LiveCodeBench-X, but a significant improvement on MultiPL-E.
 Model MultiPL-E Ag-LCB-X
-X= Lua Julia R X= Lua Julia R
-Qwen3-8B 63 53 44 11 9 9
-[…]
-Agnostics scales to larger models
-To test if improvements from Agnostics training scale with model size, we train the Qwen 3 8B model on Ag-Codeforces-X specialized to Lua, Julia and R and benchmarked it on Ag-LiveCodeBench-X and MultiPL-E (Table 3).
-The Qwen3-8B-CF-X models improve significantly on both benchmarks, improving over their 4B counterparts.
-We expect Agnostics to scale to even larger models, with appropriate computing resources.
-However, we found that Agnostics training with Ag-Codeforces-X does not improve two smaller models: Qwen 3 1.7B and Llama 3.2 3B Instruct, perhaps due to the problems being too difficult.
-Table 4: Pass@1 rates on for Qwen 3 4B models trained using MBPP. This smaller and easier dataset still results in improvements, though smaller than those in Tables 1 and 2.
-Model MultiPL-E Ag-LCB-X
-X= Lua Julia R X= Lua Julia R
-Qwen3-4B 51 61 36 11 10 10
-[…]
-Agnostics works with easier problems
-So far, all of the models that we have trained so far have used the Open-R1 Codeforces dataset.
-To show that Agnostics works with other datasets, we also train models for Julia, Lua, and R using the MBPP training set.
-(§ 3.1 describes how we prepare MBPP.) Note that the MBPP problems are trivial—see figure 3(a)—compared to the Open-R1 Codeforces problems.
-So, we cannot expect these models trained on the MBPP problems to be as good as the models we presented before.
-Nevertheless, training on MBPP still improves performance on Lua and Julia (table 4). On R, the table shows a small drop in performance on Ag-LiveCodeBench-X, but an improvement on MultiPL-E.
-Table 5: Pass@1 rates for DeepSeek Coder V1 and Phi 4 models trained using Ag-Codeforces-X. We find that Agnostics works on other model families.
-Model MultiPL-E Ag-LCB-X
-X= Lua Julia R X= Lua Julia R
-Phi4-mini-ins 40 39 34 7 7 0
-DSC-6.7B-Ins 40 54 37 8 5 7
+X= Lua Julia R Lua Julia R
+SmolLM3 3B 11 12 18 1 2 2
+Phi 4 mini ins 40 39 34 8 8 5
+DSC 6.7B Ins 40 54 37 8 5 8
 […]
 […]
-Agnostics works on other model families
-Finally, to test if Agnostics works on other models, we train both DeepSeek Coder 6.7B Instruct (Guo et al.
-2024) and Phi 4 Mini Instruct (Microsoft et al.
-2025) with Ag-Codeforces-X specialized to Julia, Lua, and R.
-Agnostics is able to improve the performance of both models on all languages, as measured by MultiPL-E and Ag-LiveCodeBench-X with one exception (Table 5).
-Phi 4 Mini Instruct scores zero on Ag-LiveCodeBench-R, and training does not improve the score.
-However, on the easier MultiPL-E benchmark, we see that the model is slightly better on simpler R problems ($34\rightarrow 35$).
-Note that DeepSeek Coder 6.7B is a relative old LLM, superseded by the much larger DeepSeekV2 and V3 models.
+[…]
+Table 4: Non-Qwen models, pass@1.
+Agnostics works on multiple model families
+To show that Agnostics works on non-Qwen models, we train SmolLM 3 (SmolLM3 Team, 2025), Phi 4 Mini Instruct (Microsoft, 2025), and DeepSeek Coder 6.7B Instruct (Guo et al., 2024) on Ag-Codeforces-X specialized to Lua, Julia, and R.
+Agnostics improves these models’ performance on all languages, as measured by MultiPL-E and Ag-LiveCodeBench-X (Table 4).
+Note that DeepSeek Coder 6.7B is a relatively old LLM, superseded by the much larger DeepSeekV2 and V3 models.
 Unlike Qwen 3, DeepSeek Coder is not trained with reinforcement learning, but is only an instruction-tuned model.
 Thus this result also shows that Agnostics can work on models that have had relatively limited post-training.
-Figure 6: We classify with an LLM what bugs appear in programs produced by Qwen 3 4B and our trained model, Qwen3-4B-CF-OCaml.
-Each radial represents a bug category, and points along the radial indicate how many programs appear to have that bug.
-Qwen 3 4B makes more fundamental programming mistakes, such as syntax errors and misusing builtins, indicating its limited grasp of OCaml.
-Our trained model makes significantly fewer such mistakes.
+Model Ag-LCB-Fortran
+Sonnet 4 Thinking (teacher) 12
+Qwen 3 4B (student) 0
+[…]
+[…]
+[…]
+Table 5: Distillation experiment results.
+Agnostics outperforms distillation
+So far we discussed training a model on its generations.
+An alternative is to distill a larger model (assuming one exists).
+As larger models do not perform very well on many low-resource programming languages, one can expect distillation to be less effective.
+We run the following experiment to verify this claim.
+Using Sonnet 4 Thinking, we synthesize Fortran solutions to Ag-Codeforces-X problems, creating a training set of 1,987 items.
+(For 13 items, Sonnet 4 (sonnet-4-20250514) with extended thinking does not produce a response within its reasoning budget.) To make sure generating the training items does not use significantly more compute compared to Agnostics training, we use at most 32K reasoning tokens, spending $\mathdollar 96$ to generate the items.
+We fine-tune Qwen 3 4B for 3 epochs (batch size 64, learning rate $2\times 10^{-5}$, cosine learning rate decay with warmup ratio $0.1$).
+Table 5 shows the resulting models reach scores far lower than the $15\%$ of Qwen3-4B-CF-Fortran (table 2).
+Agnostics outperforms rejection sampling
+We also ran a small experiment to confirm that rejection sampling would be prohibitively expensive in our case.
+In rejection sampling with supervised fine-tuning , we prompt the model to synthesize $n$ solutions to each task, reject solutions that fail tests, and fine-tune on the task-solution pairs that pass tests.
+Cassano et al.
+(2024) use this approach to get solutions for ${\approx}30\%$ of the tasks in their dataset.
+The efficiency of this approach depends on the hardness of the task and the capabilities of the model.
+In this paper, we use newer models that are marginally better at low-resource languages (based on MultiPL-E benchmark results).
+The task of Cassano et al.
+(2024) is to translate a simple, self-contained Python function from the model’s pretraining data into an equivalent function in another programming language.
+This task is significantly easier than the Ag-Codeforces-X task, which is to solve a competitive programming problem in a low-resource language without any reference code.
+Rejection sampling would be prohibitively expensive for the low-resource programming languages we consider.
+Cassano et al.
+(2024) report a $30\%$ success rate on their code translation task.
+During Agnostics training, Qwen3-4B-CF-Fortran generated a correct answer to a train split problem only $6.64\%$ of the time, generating $11400$ verified programs overall.
+We also sampled responses to the same problems from the base model of Qwen3-4B-CF-Fortran, Qwen 3 4B, taking the same amount of samples with the same generation parameters as used during training.
+The base model succeeded $0.09\%$ of the time, generating only $158$ test-passing programs.
+Figure 6: LLM labels of bugs in programs synthesized by Qwen 3 4B and our trained model, Qwen3-4B-CF-OCaml.
+A radial represents a bug class, and points along the radial show how many programs seem to have that bug.
+Qwen 3 4B makes more fundamental programming mistakes, such as syntax errors and misusing builtins, showing its limited grasp of OCaml.
+Our trained model makes far fewer such mistakes.
 We see a small increase in logic flaws: the trained model makes fewer shallow mistakes, revealing deeper issues.
 4.4 Qualitative Improvements
-In this section we take a deeper look at how training with Agnostics addresses the kinds of errors that Qwen 3 4B makes on low-resource languages.
-First, we define a taxonomy of common bugs by prompting an LLM (o3) to classify bugs in a sample of faulty programs and then lightly editing the suggestion.
-§ E has the full taxonomy and the prompt we used to develop it.
-The taxonomy spans fundamental programming errors, such as syntax errors, and more subtle mistakes such as logic flaws.
-We then sample 100 problems from Ag-LiveCodeBench-X, and take five programs produced by Qwen 3 4B and the Agnostics models trained on Open-R1 Codeforces problems.
-Recall from table 1 that these are significantly better than Qwen 3 4B.
-Using Sonnet 4, we use the taxonomy to classify the bugs in each program.
+We now take a deeper look at how training with Agnostics qualitatively improves Qwen 3 4B.
+First, we define a taxonomy of common bugs; we prompt o3 to classify bugs in a sample of faulty programs and lightly edit the suggestion.
+§ D has the full taxonomy and the prompt used to develop it.
+The taxonomy spans fundamental programming errors such as syntax errors, and subtler mistakes such as logic flaws.
+We sample 100 problems from Ag-LiveCodeBench-X, and for each take five OCaml programs produced by Qwen 3 4B and the models trained on Ag-Codeforces-X.
+Recall from table 2 that our models significantly outperform the base Qwen 3 models on our benchmark.
+With Sonnet 4, we use the taxonomy to classify the bugs in each program.
 Figure 6 shows the bug distribution for OCaml.
-We see that the base Qwen 3 4B models makes substantially more fundamental programming mistakes in OCaml.
+We see that the base Qwen 3 4B model makes many more fundamental OCaml programming mistakes.
 More programs have syntax errors (55% vs 35% after training), more programs misuse builtin functions (60% vs 32%), and so on.
-However, we observe a small increase in logic flaws (18% vs 25% after training).
-Inspecting the results, the reason for this is that when a program is full of syntax errors and hallucinated functions, it is very difficult to even determine whether or not the algorithmic approach is correct.
-Training eliminates these shallow bugs and lets deeper issues manifest.
-We see the same patterns across the other four languages (§ E).
+We do observe a small increase in logic flaws (18% vs 25% after training).
+When a program is full of syntax errors and hallucinated functions, it is hard to decide if the algorithmic approach is correct; training eliminates these shallow bugs and lets deeper issues manifest.
+Models trained for the other four programming languages show the same patterns (§ D).
 5 Conclusion
-LLMs are strongest where pre-training and post-training data are abundant, and weakest where practitioners arguably need them most: for low-resource programming languages.
-This paper proposes Agnostics, a language-agnostic post-training pipeline that removes the per-language engineering tax by verifying code purely via externally observable behavior.
-A single verifier and a short configuration are enough to adapt the same reinforcement learning setup to new languages.
-Empirically, Agnostics consistently improves small open-weight models on five low-resource languages–Lua, Julia, R, OCaml, and Fortran–without requiring language-specific test translators.
-Training Qwen-3 4B with Ag-Codeforces-X yields large gains on our new Ag-LiveCodeBench-X benchmark and on MultiPL-E, often rivaling or surpassing 16B–70B open-weight baselines.
-The method scales to larger and different model families: Qwen-3 8B benefits similarly, and we observe improvements on DeepSeek Coder 6.7B and Phi-4 Mini as well.
-Error-type analysis shows that the gains come from reducing fundamental language-usage mistakes.
+Arguably, practicioners need LLMs the most for low-resource programming languages, which need highly specialized knowledge.
+Unfortunately, this is precisely where LLMs are weak due to a lack of sufficient pre-training and post-training data.
+We propose Agnostics, a language-agnostic post-training pipeline that minimizes the per-language engineering tax by verifying code entirely via externally observable behavior.
+A short configuration file is enough to adapt the pipeline, including the reinforcement learning setup, to a new programming language.
+Empirically, Agnostics consistently improves small open-weight models on five low-resource languages—Lua, Julia, R, OCaml, and Fortran—without requiring language-specific test translators.
+Training Qwen 3 4B with Ag-Codeforces-X yields large gains on our new Ag-LiveCodeBench-X benchmark and on MultiPL-E, often rivaling or surpassing 16B–70B open-weight baselines.
+The method scales to larger and different model families: Qwen 3 8B shows similar gains to its smaller sibling, and we also observe improvements on DeepSeek Coder 6.7B, Phi 4 Mini and SmolLM3 3B.
+Error-type analysis shows our training decreases fundamental programming language mistakes.
 A practical advantage of the approach is how little per-language work it requires.
-After the framework was in place, adding OCaml and Fortran took us less than an hour each.
-We expect adoption to be just as straightforward for any language with a command-line toolchain.
-We see no reason to expect the approach to not work for models larger than 8B, although our experiments are limited by available compute.
-For scaling data, the Agnostics reformulation strategy apply to much larger problem sets as well.
-For instance, OpenCodeReasoning has $\sim$600K problems with Python solutions (Ahmad et al.
-2025); converting such corpora into language-agnostic I/O tasks would provide rich RL datasets with many target languages with minimal additional engineering.
-6 Acknowledgments
-This work was supported by Czech Science Foundation Grant No.
-23-07580X.
+After the framework was in place, extending the pipeline to support OCaml and Fortran took us less than an hour each.
+We expect adaption to be just as straightforward for any pragmatic programming language with a command-line toolchain.
+We believe the approach scales to models of arbitrary size, although our experiments are limited by available compute to at most 8B models.
+For scaling data, the Agnostics reformulation approach also applies to much larger problem sets.
+For instance, OpenCodeReasoning has $\sim$600K problems with Python solutions (Ahmad et al., 2025); converting such corpora into language-agnostic I/O tasks would provide rich RL datasets with many target languages with minimal additional engineering.
+Reproducibility Statement
+We provide references to our data, code, and artifacts at agnostics.abgru.me.
+We share our datasets and models at huggingface.co/collections/nuprl/agnostics, our Wandb training logs at wandb.ai/nuprl/Agnostics, and our training framework at github.com/nuprl/agnostics-framework.
+The existing datasets we used are publicly available and are accompanied by citations.
+We publicly release all datasets we introduce, allowing free use for research.
+In the same way, we release all code required for conducting and analyzing our experiments, including the code for dataset preparation, as well as the models we presented and Wandb records from training them.
+We state the number and range of values tried per (hyper) parameter, and outline how we chose the final values and what they are (§§ C.1 and 3.3).
+We specify the computing infrastructure (hardware and software) we used for our experiments (§ C.5).
+The released codebases specify the exact versions of all the libraries we used.
+Acknowledgments
 This work is partially supported by the U.S.
@@
 Department of Energy, Office of Science under Award Number DESC0025613.
-We used GNU parallel (Tange 2023).
+This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award ALCC-ERCAP 0038272 (m5083-2024).
+This work is supported by Czech Science Foundation Grant No. 23-07580X.
+We thank Northeastern Research Computing for support with the Northeastern University Explorer cluster.
+This work used the Delta cluster at the National Center for Supercomputing Applications (NCSA) through allocation CIS230213 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S.
+National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
+This work uses vLLM (Kwon et al., 2023), Transformers (Wolf et al., 2020), and GNU Parallel (Tange, 2023).
 Disclaimer: This report was prepared as an account of work sponsored by an agency of the United States Government.
@@
 The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
-Appendix A Efficiency of Rejection Sampling
-An alternative to reinforcement learning is to use rejection sampling with supervised fine-tuning: prompt the model to synthesize $n$ solutions to each task, reject solutions that fail tests, and fine-tune on the task-solution pairs that pass tests.
-If the model is weak at a task, which is the case with low-resource languages, then it is possible for all solutions to fail for a given task.
-Cassano et al.
-(2024) use this approach to get solutions for $\approx 30\%$ of the tasks in their dataset.
-They use $n=50$ for Lua and Julia, but increase to $n=100$ for OCaml.
-The efficiency of this approach depends on the hardness of the task and the capabilities of the model.
-In this paper, we work with newer models that are marginally better at low-resource languages (based on MultiPL-E benchmark results).
-The task in Cassano et al.
-(2024) is to translate a simple, self-contained Python function from the model’s pretraining data into an equivalent function in another programming language.
-This task is significantly easier than the Agnostics task, which is to solve a competitive programming problem in a low-resource language, without any reference code or tests.
-We can estimate the efficiency of rejection sampling from the rewards assigned in the early steps of reinforcement learning, when the policy is closest to the initial policy.
-Figure 7 shows the pass1 rates while training Qwen 3 4B on Fortran and OCaml.
-In the first 200 steps, the average pass rate for Fortran and OCaml are $1.9\%$ and $1.1\%$ respectively–considerably less than the $30\%$ success rate reported by Cassano et al.
-(2024) on their code translation task, making the rejection sampling approach prohibitively expensive for the low-resource language setting.
-Appendix B Preparing Datasets for Agnostics
+Appendix A Preparing Datasets for Agnostics
 Ag-MBPP-X
-is a dataset of Mostly Basic Programming Problems, transformed from the MBPP (Austin et al.
-2021) dataset.
+is a dataset of Mostly Basic Programming Problems, transformed from the MBPP (Austin et al., 2021) dataset.
 The source dataset contains problems asking to complete a Python function from on its signature and a docstring, where the docstring specifies what result the function should return; our dataset contains equivalent programs which read/write the same data from/to standard input/output.
 Out of 974 problems, we are able to translate 776 problems into Ag-MBPP-X (348 problems are in the sanitized subset of MBPP).
+We analyzed Ag-MBPP-X for data similarity against Ag-LiveCodeBench-X using Decon (2025) and found no data overlap.
+The analysis was done based on 5-grams with no token sampling (see Decon documentation).
 Processing the dataset with Qwen3-32B took less than 1 hour using 2 H100 GPUs.
@@
 We prepared both a train and a test split.
-The former contains 5369 problems and the latter contains 107 problems we held out from the source dataset, 7 selected manually and 100 randomly.
+The former contains 5369 problems and the latter contains 105 problems we held out from the source dataset, 5 selected manually and 100 randomly.
+The manually-selected problems were chosen to be much easier than average, to make it easier to detect if a model can solve any problems at all in a given programming language.
+In short, the 5 problems are: “output a number in binary notation”, “remove all digits from a string”, “check if the parentheses are balanced”, “parse two integers and add them”, and the following longer problem: “Petr stands in line of n people, but he doesn’t know exactly which position he occupies.
+He can say that there are no less than a people standing in front of him and no more than b people standing behind him.
+Find the number of different positions Petr can occupy.”
 The train split features randomized prompts, which we found help with generalizing the results of training on the dataset to other benchmarks.
@@
 30% of the prompts use standard Markdown headings to start different sections of the prompt, 35% use bold text instead, and the remaining 35% simply concatenate the prompt sections together.
-Half of the prompts of the final type also do not include an I/O sample in the prompt.
+Most of the prompts follow the source dataset and feature an I/O sample in the prompt, with other samples withheld as private.
+Half of the prompts of the final type do not feature any I/O sample.
 Ag-LiveCodeBench-X
-is also a dataset of competitive programming problems, created from a subset of the LiveCodeBench dataset (Jain et al.
-2024b).
+is also a dataset of competitive programming problems, created from a subset of the LiveCodeBench dataset (Jain et al., 2024a).
 LiveCodeBench 5.0 has 880 problems, of which 381 have Python starter code and test cases.
@@
 Hence we used these problems to transform LiveCodeBench into an Agnostics dataset.
-Ag-LiveCodeBench-X only has a train split, like its source dataset.
-Appendix C Agnostics Configurations
+Ag-LiveCodeBench-X only has a test split, like its source dataset.
+Appendix B Agnostics Configurations
 In this section, we list the configurations that we use for our target languages. The configuration files use YAML. The prompts for OCaml and Fortran have instructions generated by OpenAI o3.
@@
 execute: julia snippet.jl
-The R configuration (the same as figure 4, repeated here for convenience):
+The R configuration (unmodified, unlike figure 4):
 install: apt-get install -y r-cran-tidyverse
@@
 prompt: |
-Use R version 4.
-Use readLines(con = file("stdin"))‘ to
-read input from stdin. Optionally, use
-the n argument to read the first n
-lines. For example:
+Use R version 4. Use readLines(con = file("stdin"))‘ to read input from stdin. Optionally, use the n argument to read the first n lines. For example:
 ‘‘‘r
@@
 n <- as.integer(input)
-print(n) # print the first line of input
-Also, use cat to print output to stdout.
-For example:
+cat(n) # print the first line of input
+Also, use cat to print output to stdout. For example:
 ‘‘‘r
@@
 Use Fortran 90. Some tips:
-Always begin each scope with implicit none, pick explicit kinds via
-selected_*_kind, and declare proper lengths-character(len=*) is legal only for
-dummy arguments, not locals. Strings are blank-padded: call len_trim before
-iterating, and store dynamic text in deferred-length allocatables
-(character(len=:), allocatable :: s). List-directed read(*,*) arr does not
-auto-size arrays; read a count first, then allocate and read, or tokenize a
-line manually. When translating 0-based formulas (heaps, bit positions)
-remember Fortran arrays default to 1-based; if you want 0-based, declare lower
-bounds. Use real literals (2.0d0, 1.0_rk) to avoid silent integer division,
-and guard against overflow when exponentiating integers. For frequency tables,
-allocate an array or use findloc; Fortran lacks native dicts/sets, so you must
-implement search yourself. Prefer array intrinsics (sum, count, pack) over
-hand-rolled loops, and keep helper procedures inside a contains section or
-module so interfaces are explicit. return inside the main program is
-non-idiomatic; use structured blocks or stop. Never print interactive prompts
-in batch solutions; just read, compute, and write.
+Always begin each scope with implicit none, pick explicit kinds via selected_*_kind, and declare proper lengths-character(len=*) is legal only for dummy arguments, not locals.
+Strings are blank-padded: call len_trim before iterating, and store dynamic text in deferred-length allocatables (character(len=:), allocatable :: s).
+List-directed read(*,*) arr does not auto-size arrays; read a count first, then allocate and read, or tokenize a line manually.
+When translating 0-based formulas (heaps, bit positions) remember Fortran arrays default to 1-based; if you want 0-based, declare lower bounds.
+Use real literals (2.0d0, 1.0_rk) to avoid silent integer division, and guard against overflow when exponentiating integers.
+For frequency tables, allocate an array or use findloc; Fortran lacks native dicts/sets, so you must implement search yourself.
+Prefer array intrinsics (sum, count, pack) over hand-rolled loops, and keep helper procedures inside a contains section or module so interfaces are explicit.
+return inside the main program is non-idiomatic; use structured blocks or stop.
+Never print interactive prompts in batch solutions; just read, compute, and write.
 install: apt-get install -y gfortran
@@
 execute: ./snippet.out
-Appendix D Training and Results
-D.1 Choosing hyperparameters
+Appendix C Training and Results
+C.1 Choosing hyperparameters
 Before picking the hyperparameters described in § 3.3, we investigated other values by training the Qwen 3 4B model on a previous version of Ag-Codeforces-X.
@@
 We decided against it since some of the runs degraded the model’s capabilities, unlike any of the runs we did with a cosine decay schedule.
-We also compared between GRPO group sizes of 16, 32 and 64 by training the same model on Lua.
-We observed that in some runs with group size 16, the model improved significantly less than at higher group sizes.
-Since we found no significant difference between group sizes 32 and 64, we chose the smaller one due to limited resources.
-D.2 Training dynamics
+We compared between GRPO group sizes of 16, 32 and 64 by training the same model on Lua.
+In some runs with group size 16, we saw the model improved significantly less than at higher group sizes.
+We ran an experiment to compared different temperature and group size settings (table 6).
+Table 6: Hyperparameter sweep, pass@1 score.
+Model Group size Temperature Ag-LiveCodeBench-X
+Qwen3-4B-CF-Lua 32 0.7 23.00
+normal-r1 32 0.7 19.87
+normal-r2 32 0.7 21.58
+size16-r1 16 0.7 19.80
+size16-r2 16 0.7 19.40
+size16-r3 16 0.7 19.65
+size16-r4 16 0.7 20.61
+size64-r1 64 0.7 21.71
+size64-r2 64 0.7 21.21
+size64-r3 64 0.7 20.87
+temp0p2-r1 32 0.2 19.91
+temp0p2-r2 32 0.2 20.39
+temp0p2-r3 32 0.2 21.98
+temp1-r1 32 1.0 21.22
+temp1-r2 32 1.0 21.48
+temp1-r3 32 1.0 20.30
+The models trained at group size 16 had slightly lower scores compared to other models, while the ones trained at group size 64 displayed scores comparable to other models.
+However, they took significantly longer to train.
+Two group size 64 models took $\sim 20.5$h to train on average (the third one was trained on a different machine).
+In comparison, the group size 32 models trained at the same time on the same machine took $\sim 12$h on average.
+The models trained at temperatures other than the $0.7$ recommended by the Qwen team performed similarly to the other models.
+As we found no significant difference between the temperature settings and between group sizes 32 and 64, we chose the smaller group size due to limited resources, and used the recommended temperature settings.
+C.2 Training dynamics
 Figure 8: Training Qwen3-4B-CF-X, GRPO group batch pass@1.
@@
 Figure 8 shows the GRPO group batch pass@1 while training the Qwen3-4B-CF-X models.
-The scores of all the models are broadly correlated with one another, in part due to training them on the same permutation of the training data.
+The scores of all the models are broadly correlated with one another, which may at least in part be due to training them on the same permutation of the training data.
 Figures 10, 11, 12, 13 and 14 compare the GRPO group pass@1 scores with the pass@1 scores on the test split.
@@
 In most cases, we see that the train scores keep increasing until the end of the epoch, together with the test split pass@1 scores, indicating that the model keeps improving until the end of the dataset.
-D.3 Hardware and Software Used
+C.3 Reward function
+We investigated the results of partial rewards.
+We trained Qwen 3 4B—the base model of Qwen3-4B-CF-X—on Ag-Codeforces-Lua, giving it a partial reward of $0.2$ if it generated code which failed one of the tests by producing wrong output but otherwise terminated without an error.
+The full reward for a snippet passing all tests was still 1.
+Table 7 shows that the trained models score below Qwen3-4B-CF-Fortran both on Ag-LiveCodeBench-Lua and on the test split of Ag-Codeforces-Lua (counting only the full-credit reward).
+The latter scores are particularly far lower, clearly showing that the models learned to abuse the partial-credit reward.
+During training, we saw the models focus on the partial reward.
+The average result from the partial reward component was clearly increasing more quickly than the result from the full reward component.
+In the training generations we inspected, the models also often claimed to generate a “draft” answer and produced a program which ignored the problem in the prompt, for instance by only printing a hard-coded string such as “0”.
+Table 7: Partial reward experiment, pass@1 rates.
+Model Ag-LiveCodeBench-X Ag-Codeforces-X (test split)
+Qwen3-4B-CF-Lua 23.00 24.76
+partial-r1 18.57 13.33
+partial-r2 20.16 15.62
+C.4 Cross-Programming-Language Negative Transfer
+To demonstrate that Agnostics training does not lower performance on different programming languages, we evaluated the models we trained on variants of Ag-LiveCodeBench-X (table 8).
+Table 8: Cross-PL evaluation, pass@1 rates.
+Model Ag-LiveCodeBench-X
+X= Python Lua Julia R
+Qwen 3 4B 34.34 11.00 10.00 10.00
+[…]
+[…]
+[…]
+Qwen 3 8B 33.51 11.00 9.00 9.00
+[…]
+[…]
+[…]
+DSC 6.7B Ins 16.86 7.89 4.79 7.97
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+C.5 Hardware and Software Used
 We used three machines while working on this paper: B, R1 and R2.
@@
 When developing the Agnostics framework, we used the following major Python libraries: ray v2.46.0, torch v2.6.0, transformers v4.54.1, vllm v0.8.5.post1, datasets v3.4.1, wandb 0.19.11.
-Appendix E Bug Taxonomy
-E.1 Prompt for Generating Taxonomy
+Appendix D Bug Taxonomy
+D.1 Prompt for Generating Taxonomy
 We used the following instructions to generate the bug taxonomy, followed by a list of faulty R programs.
@@
 Supply a short explanation for each theme in your taxonomy.
-The prompt produced a taxonomy of 11 bug categories. We edited these categories and selected 7 categories relevant to us, shown in §E.2.
-E.2 Bug Taxonomy Used For Analysis
+The prompt produced a taxonomy of 11 bug categories. We edited these categories and selected 7 categories relevant to us, shown in §D.2.
+D.2 Bug Taxonomy Used For Analysis
 The following categories represent the prevalent themes of programming errors we use in our analysis of bugs in model-generated code.
@@
     Output Formatting and Presentation Errors: Correct computational results, but incorrect due to formatting issues (e.g. missing newlines/spaces or output spec violations).
-E.3 Radar Charts for All Programming Languages
+D.3 Radar Charts for All Programming Languages
 Figures 15, 16, 17, 18, 19 show the error theme charts for all the programming languages we trained a model on.
@@
 Figure 19: Radar chart of Fortran error themes for Qwen3-4B and Qwen3-4B-CF-Fortran.
-Appendix F Reproducibility Checklist
-This paper:
-  • Includes a conceptual outline and/or pseudocode description of AI methods introduced: yes
-  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results: yes
-  • Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper: yes
-Does this paper make theoretical contributions? No.
-Does this paper rely on one or more datasets? Yes.
-  • A motivation is given for why the experiments are conducted on the selected datasets: yes.
-  • All novel datasets introduced in this paper are included in a data appendix: no.
-  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: yes.
-  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations: yes.
-  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available: yes.
-  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisfying: N/A.
-Does this paper include computational experiments? Yes.
-  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting: yes.
-  • Any code required for pre-processing data is included in the appendix: no.
-  • All source code required for conducting and analyzing the experiments is included in a code appendix: no.
-  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: yes.
-  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from: no.
-  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results: yes.
-  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks: yes.
-  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics: yes.
-  • This paper states the number of algorithm runs used to compute each reported result: yes.
-  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information: no.
-  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank): no.
-  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments: yes.